### v1

# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

Abstract
In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras. In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model.Our model serially connects a camera Next-Best-View (NBV) policy with a gripper Next-Best Pose (NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning. This approach allows the agent to adjust a third-person camera to actively observe the environment based on the task goal, and subsequently infer the appropriate manipulation actions. We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench. The results demonstrate that our model consistently outperforms baseline algorithms, showcasing its effectiveness in handling visual constraints in manipulation tasks. Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
Index Terms:
Reinforcement learning , Robotic manipulation, Active vision, Next-best-view, Sensorimoter.
I Introduction
Despite the reinforcement learning based models for robotic manipulation can learn to directly infer effective actions from camera-provided observation images through interaction with the environment, exhibiting strong scene perception and reasoning capabilities, they still heavily rely on cameras positioned based on prior knowledge, such as third-person fixed cameras and first-person wrist-mounted camera[1][2][3]. While these fixed third-person cameras and wrist-mounted cameras help reduce the sample variance of observations during interaction[3][4], they often struggle with providing sufficient visual information when dealing with manipulation scenarios under occlusions and limited observations due to unadjustable viewpoint positions or fixed camera-gripper configuration.
Humans can independently adjust their head position and eye fixation based on their understanding of task goals and the current environment state to obtain the most advantageous visual information for making interactive action decisions[5]. Inspired by this, recent works in active vision-action joint learning[3][6][7][8] aim to mimic human capabilities by enabling agents to actively adjust their sensory “eye” viewpoints and interact with the world using their motor “hands.”
Similar to humans, the active vision-action model exists in two categories: synchronous and asynchronous[10]. The synchronous vision-action model makes real-time decisions on the next visual sensor and motor actions based on current observation information[7][8][11][12]. This setup requires the model to effectively coordinate sensor motor actions in the next moment and to infer motor actions even when sufficient observation information may not be available which impose higher demands on the model’s dynamic modeling capabilities, usually requiring more interaction steps and making it difficult to apply to manipulation tasks requiring large viewpoint changes.
Although humans are capable of making synchronous decisions, hand movements and changes in viewpoint typically do not occur simultaneously; instead, we generally select the optimal eye viewpoint before executing an interaction. Similarly, the asynchronous vision-action model separates visual sensing and motor actions over time, allowing the vision policy to focus less on coordinating both actions simultaneously when determining the observation viewpoint. Instead, the model can focus more on task-specific manipulations and distribute the sensor-motor coordination throughout the entire task cycle. Although such models are currently more often applied to object grasping tasks in occluded scenes[13][14], we believe that the asynchronous approach may be more suitable for common sequential manipulation tasks.
In this paper, we further explore the problem of robotic manipulation under limited observation and introduce a task-driven asynchronous active vision-action model. The model is designed as a sequential dual-agent structure, as illustrated in Fig.LABEL:fig:head_image. It consists of two agents: a Next-Best-View (NBV) agent for inferring optimal camera viewpoints, and a Next-Best-Pose (NBP) agent for determining the gripper’s action based on the observations from the NBV-inferred viewpoint. By alternating between sensor and motor action inference in each episode, our model enables active perception of the environment and effective interaction with the environment to achieve the manipulation task objectives.
We train our model using a few-shot reinforcement learning approach[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the challenge posed by the expanded observation-action space resulting from the active viewpoint, we employ a viewpoint-centric voxel alignment technique, which rotates the scene voxel to align with the camera viewpoint. To further stabilize the training process, we incorporate a combined task-agnostic auxiliary reward which provides sufficiently task-aligned learning signals throughout the training process, encouraging the agent to interact effectively with the environment and select viewpoints that offer richer visual information for more informed gripper action decisions. We conduct comparison experiments and analyses on 8 manipulation tasks with limited effective viewpoints in RLbench[9]. The results demonstrate our model’s effectiveness in selecting task-relevant viewpoints and its robustness in inferring manipulation actions based on the current state and task requirements. The main contributions of this paper are as follows:
  • We propose a task-driven asynchronous active vision-action model that uses a novel dual-agent system to achieve effective robotic manipulation under limited observation.
  • We introduce methodological innovations, including viewpoint-aware demo augmentation, viewpoint-centric voxel alignment, and a task-agnostic auxiliary reward strategy, to enhance model performance, learning efficiency, and training stability.
  • Comprehensive experiments on public RLbench tasks validate the effectiveness and robustness of our approach.
II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12][15], object searching[16, 17, 18, 19] and 3D reconstruction[20]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7][13][14][21][22] and for detecting fruit nodes in automated harvesting[23, 24, 25]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
3D Voxel-based Manipulation. 3D voxels are highly effective in representing spatial structures[26], making them widely used in domains such as medical imaging[27][28] and autonomous driving[29][30]. Recently, they have also been adopted in robotic manipulation models for visual representation. For instance, James et al.[2] transformed original RGB-D images into 11-dimensional voxel grids and developed a 3DUnet-based reinforcement learning model to extract scene features and infer manipulation actions. Similarly, Shridhar et al.[31] used discrete 3D voxel images as input, employing the Perceiver network[32] as a backbone for end-to-end behavioral cloning, trained on demonstration trajectories. Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
Figure 2: Comparison of regular RL and asynchronous active vision-action decision processes. We introduce an intermediate time step $t^{\prime}$ between $t$ and $t+1$, dividing the interaction into NBV policy for viewpoint selection and the NBP policy for action selection. The task reward $r_{\text{t'}}$ for state $\mathbf{s}$ is derived from the reward $r_{\text{t+1}}$ after the NBP interaction, which enables sensor-motor joint training through shared task rewards.
Attention-driven Robotic Manipulation. The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs, improving both training efficiency and model performance[33]. ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy 2D regions via DQN, while C2FARM predicts regions of interest (ROIs) in 3D space. Liang et al. [34] pre-train a scene saliency attention encoder offline using human-annotated saliency maps and then train the reinforcement learning operation agent online based on the scene features output by this encoder. However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints. Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
III background
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optmal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
Problem Formulation. We divide one asynchronous vision-action time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy from $t^{\prime}\sim t+1$. Therefore the asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},\mathcal{O^{\prime}},P^{\prime},R,\gamma>$ As shown in Fig.2, $\mathcal{O}$ and $\mathcal{O^{\prime}}$ represent the observation spaces of the agent at times $t$ and $t^{\prime}$, respectively. The NBV policy $\pi_{\text{v}}$ determines the camera viewpoint action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, and obtains a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ through transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{g}}$ then determines the gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$ based on the new observation $\mathbf{o}_{t^{\prime}}$. Through the transition probability $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$, scene observation $\mathbf{o}_{t+1}$ can be obtained by camera at time $t+1$, and the environment provides the task reward $r_{t+1}\in R$. Our goal is to jointly learn the policies $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{g}}$ that maximize the expected sensor-motor joint reward for the given manipulation task, which is $\pi^{*}_{\text{v}},\pi^{*}_{\text{g}}=\arg\underset{\pi_{\text{v}},\pi_{\text{p}}}{\max}\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
Observation and Action Space We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\mathcal{O^{\prime}}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid voxelized from the scene RGB-D image captured by the camera, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the raw RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, which is then voxelized through function $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center coordinates. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and a occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
Following common settings[14][21][22], we restrict the camera viewpoint’s movement to the upper hemisphere at a fixed radius of $r^{\text{v}}=1.2\text{m}$ centered on the table, illustrated in Fig.8. The camera’s optical axis is assumed to always be directed towards the table center thus the viewpoint position can be described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. We than define the NBV policy’s action as $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}\in\mathcal{A}^{\text{c}}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy’s action is defined as $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}\in\mathcal{A}^{\text{g}}$, where $\mathbf{k}^{\text{g}}$ represents the discrete 6-DoF gripper pose index, and $d$ is the goal distance between the parallel gripper tips.
IV Method
In this section, we provide a detailed explanation of the fundamental structure of our active vision-action model, along with three innovative strategies designed to enhance both model performance and learning efficiency.
IV-A Model Structure
Fig3 illustrates the basic pipline of our model. At time $t$, the voxelized image observation and proprioceptive state undergo a coordinate transformation for viewpoint alignment (Section IV-B), resulting in the processed observation $\mathbf{o}_{t}$. The NBV policy network, parameterized by $\Theta$, uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, representing the viewpoint that would best improve task success. The camera is then adjusted to the goal viewpoint $\mathbf{v}^{*}$, which is derived from the selected $\mathbf{k}^{\text{v}}$.
$\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
Based on the predicted ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the raw image observation from the viewpoint is cropped and aligned to generate the corresponding ROI 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$. This enables us to construct the observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Using the NBP policy network parameterized by $\Phi$, we infer the gripper closure action $d$ and the discrete gripper pose index $\mathbf{k}^{\text{g}}$ which is then transformed into the goal pose $\mathbf{g}^{*}$. Finally, a complete sensor-motor interaction step is achieved after gripper exectuation and rewarded by $r_{t+1}$.
$\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
We denote the sub-transition for the NBV policy as $\text{T}_{t}^{\text{v}}=<\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}}>$ and for the NBP policy as $\text{T}_{t}^{\text{p}}=<\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1}>$. Since the NBV and NBP policies share the same reward $r_{t+1}$, a complete interaction transition is denoted as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$. Upon completing an interaction episode, the collected transition is stored in the replay buffer together with the demonstration transitions. During model updates, both the NBV and NBP policies are updated using the same reward, thereby achieving sensor-motor joint learning. We employ a 3D-CNN based U-Net[35] as the image encoder for both policy networks, and three-layer fully connected network (FCN) for decoding the viewpoint and gripper actions.
[…]
IV-B Viewpoint-centric Voxel Alignment
Viewpoint movement expands the image observation space, increasing the variance of interaction samples[3]. Specifically, images captured from a static viewpoint are primarily influenced by the scene state, following the mapping $f^{\text{static}}:\mathcal{S}\rightarrow\mathcal{O}^{\text{v}}$. In contrast, active viewpoint movement introduces an additional influence from the viewpoint itself, resulting in the mapping $f^{\text{active}}:\mathcal{S}\times\mathcal{V}\rightarrow\mathcal{O}^{\text{v}}$. Since the 3D-CNN lacks spatial rotation invariance, voxelizing the scene using the world coordinate system as the reference frame can cause features of similar structures from different viewpoints to appear distant in the feature space after encoding, as shown in Fig.4. This misalignment negatively impacts the overall training efficiency of the model.
[…]
$\mathbf{o}^{\text{v}}=\mathbf{V}\big{(}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}}),\mathbf{e},\mathbf{c}\big{)}$ (3)
[…]
$\mathbf{v}^{*},\mathbf{f}^{*},\mathbf{g}^{*}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}^{*}_{\text{L}},\mathbf{f}^{*}_{\text{L}},\mathbf{g}^{*}_{\text{L}})$ (4)
IV-C Viewpoint-aware Demo Augmentation
Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations (represented by hatched circles) between both viewpoint key frames (solid blue circles) and gripper key frames (solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
To provide early guidance and reduce the need for extensive exploration, we manually alternate moving the camera and gripper to generate and collect a small number of demonstration trajectories as high-quality examples. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of “observation-action” pairs collected at a fixed sampling rate. Existing demonstration augmentation methods[1][2][31] typically involve two main processes: key frame discovery and observation sampling. In the key frame discovery process, a limited set of key frame indices is selected from a trajectory $\tau_{d}$, and transition samples are constructed based on the corresponding “observation-action” pairs. The observation sampling process then generates additional transition samples by resampling the initial observations between the identified key frames. However, existing methods do not account for viewpoint movement. Our method incorporates key frame discovery and observation resampling, taking into account viewpoint movement, enabling viewpoint-aware demonstration augmentation. Specifically, we define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, which retrieves the key frames of both viewpoint and gripper movements in the demonstration trajectory based on predefined rules.
The discovery of gripper key frames $k_{j}^{\text{g}}$ follows the method described in [1]. We define key viewpoints as the first stable viewpoint between every two gripper key frames, corresponding to the frame index $k_{j}^{\text{c}}$. The relationship between key frame indices satisfies $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, after the key frame discovery process, we construct the raw transition samples $\text{T}_{j}$ formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{<\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}>,<\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented frame indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ &=\{<\mathbf{o}_{(k_{j-1}^{\text{g}})^{\text{aug}}},\mathbf{a}_{(k_{j-1}^{\text{g}})^{\text{aug}}}^{\text{c}},\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}}>,\\ &<\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{a}^{\text{g}}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (6)
Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. We encourage the model to select non-empty ROIs (green dashed boxes) and maximize the number of occluded voxels within the ROI that can be observed from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$ compared to the original viewpoint $\mathbf{v}_{t}$. This strategy promotes active interaction with the environment and enables the model to capture more structural information within the ROI.
IV-D Task-agnostic Auxiliary Reward
We use the task completion signal as a sparse task reward to avoid designing targeted rewards for different tasks; however, sparse reward only provides positive feedback to the model at the end of an episode, making it difficult to efficiently guide the model in sensor-motor joint learning. To address this, we introduce two task-agnostic auxiliary rewards that provide real-time feedbacks to the training process.
To encourage effective interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, reward $r_{\text{i}}$ provides a positive signal when the NBV policy infers an ROI in a reachable and non-empty area of the scene, and penalizes actions that select an ROI in an empty or unreachable area due to distance or trajectory planning failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI not reachable and empty}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward , denoted as $r_{\text{e}}$, inspired by [24]. The core idea is to reduce the occupancy entropy within the ROI after adjusting the viewpoint, as illustrated in Fig.6. Specifically, assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel is calculated using the Shannon entropy function, as shown in Eq.(8), where $p_{\text{o}}(i)$ represents the voxel’s occupancy probability. If a voxel is occluded, its occupancy probability is set to $p_{\text{o}}=0.5$.
$e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
The total occupancy entropy change of the ROI from time $t$ to $t^{\prime}$ is expressed as:
$\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
We define the reward $r_{\text{e}}$ as Eq.(10), where the coefficient $k$ is empirically set to 1.0.
$r_{\text{e}}=k\cdot\Delta E,k=1.0$ (10)
The auxiliary reward $r_{\text{i}}$ necessitates the cooperation of both the NBV and NBP policies, whereas $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, while the reward for updating the NBV agent is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. It is important to note that the introduced rewards are designed to be task-agnostic, which ensures that the rewards can generalize across different tasks without the need for manual tuning.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 30 demonstration trajectories for each task as a benchmark. Based on this, we calculate the relative task occlusion rates for the front, overhead, and wrist viewpoints in the demonstration trajectories of each task.
V Experiment Setup
V-A Task Setting
We trained and evaluated our model on RLBench [9], a public benchmark designed for reinforcement learning manipulation algorithms. When selecting evaluation tasks, we considered two primary criteria: 1) Single task: each task must have at least one fixed viewpoint with a high level of visual occlusion. 2) Task range: no single fixed or wrist-mounted camera can provide a low occlusion rate for all tasks. Since the region of interest (ROI) the robot focuses on may change during different manipulation phases, and there may be no clear manipulation objects (e.g., when the gripper is transferring in midair), it is challenging to use an image segmentor to locate objects and define the task occlusion rate as in previous methods[14][24]. To address this, we define the Task Occlusion Rate (TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(11), Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across different tasks, we assume that the collected demo trajectories contain the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint (oracle).
$\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (11)
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under common overhead, front , wrist-mounted cameras as well as the human-demoed active camera, based on $100$ demo trajectories. The results show that, due to task characteristics and initial scene randomization, both fixed and wrist viewpoints experience varying degrees of occlusion across the 8 tasks. In contrast, the active camera demonstrates the potential to achieve lower task occlusion.
V-B Baselines
We compaired our model against two categories baseline models: 1) visual manipulaiton model that utilze common used static third-person cameras and first-person wrist-mounted camera; 2) Active third-person camera based manipulation models.
For the first category, we compared our active vision-action model against the recent visual-based C2FARM [2] model with static front, overhead camera as well as the wrist-mounted camera settings. This comparison aims to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
For the second category, using the same third-person active camera setting, we firstly compared our model against the Behavior Cloning (BC) model. Through this comparison, we seek to investigate whether our model not only mimics the given limited demonstrations but also learns new sensor-motor skills through interaction with the environment. Additionally, we tested a Random viewpoint model, following [3], which shares the same architecture as ours but with the learning of the NBV policy network disabled (learning rate set to 0). This setup allows us to evaluate the importance of the active viewpoint selection in the NBV policy and to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random viewpoint choices.
V-C Training Setting
We first collected 30 demo trajectories for each task and stored them in the replay buffer as prior samples or the training dataset, and all models were trained from scratch.
The implementation of C2FARM was based on their official release and trained using sparse task rewards. The Behavior Cloning model was trained with the Mean Squared Error (MSE) loss, while the Random model utilized the same reward structure as our model, as described in SectionIV-D. All models were trained using the Adam optimizer with an initial learning rate of $0.0005$, with no learning rate decay applied during training. Each model was trained for 40,000 updates with a batch size of 128, and all experiments were conducted on a machine equipped with 4 NVIDIA GTX 4090 GPUs.
TABLE I: Task success rate: Our model outperform baselines on most tasks and on average . Mean and std of 100 tests are reported.
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch put rubbish in bin Take Money Out Safe Unplug Charger On Average
C2FARM(wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
C2FARM(overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
C2FARM(front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
Random(active) 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.29±0.02 0.00±0.00 0.00±0.00 0.04±0.09
BC(active) 0.00±0.00 0.00±0.00 0.02±0.01 0.19±0.03 0.51±0.02 0.00±0.00 0.38±0.03 0.02±0.01 0.14±0.19
Ours(active) 0.81±0.02 0.56±0.04 0.73±0.04 0.50±0.03 0.85±0.03 0.96±0.01 0.78±0.02 0.92±0.01 0.76±0.15
TABLE II: Task episode length. The maximum length of each task episode is limited to 10 steps. Mean and std of 100 tests are reported.
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch put rubbish in bin Take Money Out Safe Unplug Charger On Average
C2FARM(wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
C2FARM(overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
C2FARM(front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
Random(active) 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 7.76±0.27 10.00±0.00 10.00±0.00 9.72±0.75
BC(active) 10.00±0.00 10.00±0.00 9.93±0.05 8.68±0.24 6.74±0.28 10.00±0.00 7.64±0.27 9.90±0.07 9.11±1.22
Ours(active) 5.35±0.08 6.43±0.28 5.15±0.26 6.65±0.19 3.89±0.28 3.33±0.09 4.30±0.17 3.91±0.07 4.88±1.16
VI Result
Main Performance. We first analyzed the average task success rate and the average episode length across different algorithms, which provide a macro-level assessment of the models’ proficiency in learning sensor-motor coordination.
Table I presents the average success rate for tasks across 100 testing episodes under 3 evaluation seeds. Our model achieved the highest average success rate across the eight tasks, outperforming the second-best by 21%. It also exhibited relatively high manipulation performance in scenarios with significant occlusion from fixed viewpoints, such as meat_on_grill, meat_off_grill and unplug_charger.
Similarly, Table II shows that the average episode length for tasks follows the same trend as the success rate. Our model achieved the shortest average manipulation steps across all tasks, outperforming the second-best model by 28.2%. Furthermore, it demonstrated the lowest variance in manipulation steps, highlighting the model’s high proficiency and consistency after training.
Moreover, we observed that in tasks where the occlusion rate is low or the scene object positions are relatively predictable, fixed and wrist-mounted cameras benefit from smaller variance in observation samples, leading to enhanced skill acquisition. For example, in tasks such as basketball_in_hoop, press_switch, put_rubbish_in_bin, the C2FARM model achieved comparable success rates from both front and overhead views, occasionally even surpassing our model. However, it is important to note that when these fixed cameras fail to capture adequate observations, performance tends to decline sharply, leading to higher performance variance across the 8 tasks.
In contrast, our model is less sensitive to visual occlusion and consistently maintains strong manipulation performance. Under the active viewpoint setting, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), its effectiveness was constrained by the limited number of training samples, which restricted its capacity to learn interaction skills. As a result, it frequently failed to complete the assigned tasks, achieving an average success rate of only 14%. Meanwhile, the random viewpoint model performed poorly across all 8 tasks, exhibiting the weakest manipulation performance, indicating its inability to effectively learn sensor-motor coordination skills.
Figure 8: Viewpoint distribution overview: Red circles indicate the distribution of manually set viewpoints across 30 demo trajectories, while blue triangles represent the distribution of viewpoint positions during successful episodes of our model over 100 test episodes. The model’s inferred viewpoints do not strictly mimic the demonstration trajectories but instead exhibit a degree of refinement and generalization.
Figure 9: Partial rollout cases of our model. The figure presents the image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the associated gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
TABLE III: A comparison of information gain to baselines.
      Models              ASSIG       (avg/max)              AIIG       (avg/max)
      C2FARM(wrist)       0.0/0.0       N/A
      C2FARM(overhead)       0.0/0.0       0.0/0.0
      C2FARM(front)       0.0/0.0       0.0/0.0
      Random(active)       0.0/5.1%       0.0/5.2%
      BC(active)       0.7%/2.3%       5.6%/18.1%
      Ours(active)       3.5%/6.9%       6.8%/25.1%
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models during the operation process using two metrics: Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG). These metrics measure the models’ ability to acquire observation information through multi-step and single-step interactions, respectively. Specifically, we define ASSIG as the average reduction in entropy within the ROI before and after a single viewpoint movement.Similarly, the initial information gain is defined as the average reduction in entropy of the ROI after a viewpoint movement, compared to the initial viewpoint. These two metrics are given by:
$\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (12)
$\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (13)
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the number of steps required to complete episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. The statistical results in Table III show that our model significantly improves the observation of the ROI region by independently adjusting camera viewpoints, outperforming the second-best BC model by 2.8% and 1.2% in the ASSIG and AIIG scores, respectively. In contrast, static front and overhead cameras, which are unable to move, fail to achieve either single-step or initial information gain through viewpoint shifts. Similarly, the wrist-mounted camera, being unable to move independently of the gripper, is limited to inferring the next gripper actions based on the current viewpoint. This results in a zero ASSIG score and makes it unsuitable for AIIG computation.
Fig.8 illustrates the distribution of viewpoints learned by our model alongside the key viewpoints distribution from 30 demo trajectories. Notably, the viewpoints learned by our model do not strictly follow the patterns observed in the demonstration trajectories. For tasks such as put_rubbish_in_bin, press_switch, unplug_charger, and meat_off_grill, the learned viewpoints are more concentrated. In contrast, for tasks like meat_on_grill, open_drawer, and take_money_out_safe, the model explores a broader range of viewpoints. This variation indicates that our model has effectively learned to select task-relevant viewpoints through interactive learning, rather than merely replicating the active perception patterns from the demonstration data. Furthermore, Fig. 8 presents camera observations at keyframe moments of partial task episodes, showcasing our model’s capacity for active visual perception during task execution.
Figure 10: Comparison of success, timeout, and execution failure rate across models.
Environmental Modeling and Interactive Capabilities. To further evaluate the environment modeling and interaction capabilities of various models, we recorded the success rate, timeout rate, and execution failure rate over 100 test episodes for each model, as illustrated in Fig.10. The C2FARM model showed a high likelihood (up to 42%) of episodes terminating prematurely due to failures in low-level trajectory execution. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. While execution failures accounted for only 4% of the BC model’s episodes, as many as 82% of the episodes failed due to timeouts. However, our model showed only a 13% rate of episodes ending prematurely due to trajectory execution failures and a mere 9% failure rate due to timeouts. These results suggest that our model effectively integrates low-level motion execution considerations when inferring manipulation actions, resulting in more robust performance across the test episodes.
Additionally, as shown in Table IV, we analyzed the average steps to first interaction with scene objects and the average number of non-interactive steps during testing for each model. Our findings indicate that the behavior cloning exhibited a relatively high number of non-interactive steps, which aligns with the high timeout rate shown in Fig.10. This pattern suggests that behavior cloning struggles to learn precise interactive manipulations from a limited number of demonstration trajectories, leading to frequent non-interactive steps and episode timeouts. In contrast, our model consistently initiates interaction with the environment more quickly and significantly reduces the time spent on non-interactive actions. These results demonstrate that, after training, our model achieves superior accuracy in modeling environmental dynamics and understanding structural relationships.
TABLE IV: First interaction step and non-interaction step comparison.
Models First Interaction Step Non-interaction Step
C2FARM(wrist) 3.21±1.82 4.86±2.70
C2FARM(overhead) 2.97±1.70 4.41±2.77
C2FARM(front) 3.41±2.28 4.17±2.74
Random(active) 9.07±1.47 9.49±1.07
BC(active) 6.31±2.46 7.54±2.15
Ours(active) 1.64±0.61 2.15±1.03
Ablation Study In this section, we investigate the impact of three key strategies on manipulation performance and various evaluation metrics: (1) viewpoint-centric voxel alignment (Section IV-B), (2) viewpoint-aware demon augmentation (Section IV-C), and (3) task-agnostic auxiliary rewards (Section IV-D). To assess the individual and combined contributions of these strategies, we conducted separate ablation tests for each, followed by a comprehensive test removing all three strategies.
As shown in Table V, each of the three strategies contributed to notable improvements in the success rate of our proposed model. Among the three, voxel alignment and viewpoint-aware demonstration augmentation had the most substantial impact, with the success rate dropping by approximately 40% when either of these strategies was ablated, highlighting their critical roles. The auxiliary reward strategy, while less significant than the other two, still provided approximately a 6% improvement, demonstrating its contribution to overall performance.
Beyond success rate improvements, all three strategies contributed to reductions in episode length, overtime ratio, planning failure rate, and the average number of non-interactive steps per episode. This demonstrates their broader importance to the algorithm’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
When all three strategies were ablated, the performance across the eight tasks showed the most severe decline, with the average success rate dropping to just 22%, even lower than the static-viewpoint baseline. This highlights the synergistic benefits of these three strategies, as their combination significantly boosts the model’s overall performance.
TABLE V: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards
Models Success Rate Episode Length Overtime Rate Planning Failed Rate
Ours(w/o align) 0.34±0.41 7.61±2.94 0.14±0.12 0.53±0.39
Ours(w/o aug) 0.31±0.37 7.99±2.44 0.15±0.17 0.55±0.37
Ours(w/o aux) 0.70±0.19 5.52±1.51 0.15±0.10 0.16±0.12
Ours (w/o align+aug+aux) 0.22±0.34 8.46±2.41 0.22±0.22 0.57±0.34
Ours 0.76±0.15 4.88±1.16 0.10±0.06 0.15±0.11
Models First Interaction Non-interaction AIIG ASSIG
Ours(w/o align) 4.33±2.80 5.81±3.30 0.046±0.05 0.082±0.07
Ours(w/o aug) 4.49±3.34 5.74±3.58 0.014±0.05 0.020±0.03
Ours(w/o aux) 1.85±0.96 2.38±1.25 0.067±0.08 0.028±0.02
Ours (w/o align+aug+aux) 5.74±3.37 6.72±3.28 0.036±0.07 0.083±0.06
Ours 1.64±0.61 2.15±1.03 0.068±0.09 0.035±0.03
VII Conclusion and Feature Work
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal observation viewpoint accroding to the task goal and then determines the corresponding manipulation action based on the updated viewpoint’s obsevation. To enhance performance and stability, our model integrates three strategies: voxel alignment for spatial consistency, demonstration augmentation for better learning, and auxiliary rewards for real-time feedback. Together, these strategies help the model handle occlusions and improve sensor-motor coordination. Experimental results show that, compared to passive observation-based model, our model significantly improves manipulation performance through actively observation. It further learns sensor-motor coordination through interaction with the environment, rather than being restricted to a limited set of demonstration trajectories. Despite these promising results, the model’s performance can still be affected by complex occlusions and dynamic environments. Future research could explore incorporating real-time visual feedback to better handle dynamic changes in the environment.

### v2

@@
 In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras.
-In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model.Our model serially connects a camera Next-Best-View (NBV) policy with a gripper Next-Best Pose (NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning.
-This approach allows the agent to adjust a third-person camera to actively observe the environment based on the task goal, and subsequently infer the appropriate manipulation actions.
+In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model.
+Our model serially connects a camera Next-Best-View(NBV) policy with a gripper Next-Best-Pose(NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning.
+This approach enables the agent to reposition a third-person camera to actively observe the environment based on the task goal, and subsequently determine the appropriate manipulation actions.
 We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench.
@@
 Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
-Index Terms:
-Reinforcement learning , Robotic manipulation, Active vision, Next-best-view, Sensorimoter.
 I INTRODUCTION
-Despite the reinforcement learning based models for robotic manipulation can learn to directly infer effective actions from camera-provided observation images through interaction with the environment, exhibiting strong scene perception and reasoning capabilities, they still heavily rely on cameras positioned based on prior knowledge, such as third-person fixed cameras and first-person wrist-mounted camera[1][2][3].
-While these fixed third-person cameras and wrist-mounted cameras help reduce the sample variance of observations during interaction[3][4], they often struggle with providing sufficient visual information when dealing with manipulation scenarios under occlusions and limited observations due to unadjustable viewpoint positions or fixed camera-gripper configuration.
-Humans can independently adjust their head position and eye fixation based on their understanding of task goals and the current environment state to obtain the most advantageous visual information for making interactive action decisions[5].
-Inspired by this, recent works in active vision-action joint learning[3][6][7][8] aim to mimic human capabilities by enabling agents to actively adjust their sensory “eye” viewpoints and interact with the world using their motor “hands.”
-Similar to humans, the active vision-action model exists in two categories: synchronous and asynchronous[10].
-The synchronous vision-action model makes real-time decisions on the next visual sensor and motor actions based on current observation information[7][8][11][12].
-This setup requires the model to effectively coordinate sensor motor actions in the next moment and to infer motor actions even when sufficient observation information may not be available which impose higher demands on the model’s dynamic modeling capabilities, usually requiring more interaction steps and making it difficult to apply to manipulation tasks requiring large viewpoint changes.
-Although humans are capable of making synchronous decisions, hand movements and changes in viewpoint typically do not occur simultaneously; instead, we generally select the optimal eye viewpoint before executing an interaction.
-Similarly, the asynchronous vision-action model separates visual sensing and motor actions over time, allowing the vision policy to focus less on coordinating both actions simultaneously when determining the observation viewpoint.
-Instead, the model can focus more on task-specific manipulations and distribute the sensor-motor coordination throughout the entire task cycle.
-Although such models are currently more often applied to object grasping tasks in occluded scenes[13][14], we believe that the asynchronous approach may be more suitable for common sequential manipulation tasks.
-In this paper, we further explore the problem of robotic manipulation under limited observation and introduce a task-driven asynchronous active vision-action model.
-The model is designed as a sequential dual-agent structure, as illustrated in Fig.LABEL:fig:head_image.
-It consists of two agents: a Next-Best-View (NBV) agent for inferring optimal camera viewpoints, and a Next-Best-Pose (NBP) agent for determining the gripper’s action based on the observations from the NBV-inferred viewpoint.
-By alternating between sensor and motor action inference in each episode, our model enables active perception of the environment and effective interaction with the environment to achieve the manipulation task objectives.
-We train our model using a few-shot reinforcement learning approach[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories.
-To address the challenge posed by the expanded observation-action space resulting from the active viewpoint, we employ a viewpoint-centric voxel alignment technique, which rotates the scene voxel to align with the camera viewpoint.
-To further stabilize the training process, we incorporate a combined task-agnostic auxiliary reward which provides sufficiently task-aligned learning signals throughout the training process, encouraging the agent to interact effectively with the environment and select viewpoints that offer richer visual information for more informed gripper action decisions.
-We conduct comparison experiments and analyses on 8 manipulation tasks with limited effective viewpoints in RLbench[9].
-The results demonstrate our model’s effectiveness in selecting task-relevant viewpoints and its robustness in inferring manipulation actions based on the current state and task requirements.
+Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3].
+While these cameras reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
+However, humans can adjust their head and eye positions based on task goals.
+This helps them obtain advantageous visual information for action decision [5].
+Inspired by this, recent active vision-action models enable agents to actively adjust their sensory viewpoints while interacting with the environment[3, 6, 7, 8], mimicking human capabilities.
+Active vision-action models, like human perception, can be categorized as synchronous or asynchronous[10].
+Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 11, 12].
+However, the need for simultaneous coordination of sensor-motor actions places high demands on the model’s dynamic modeling capabilities and may limits its ability to handle tasks requiring large viewpoint shifts.
+In contrast, asynchronous models decouple vision sensing and motor actions over time, allowing the agent to prioritize observation before interaction and then focus on inferring actions based on the selected viewpoint, thereby distributing sensor-motor coordination across the entire task episode.
+Although such models are commonly used for object grasping in occluded scenes[13, 14], we believe that the asynchronous approach may also be suitable for sequential manipulation tasks.
+In this paper, we introduce a task-driven asynchronous active vision-action model for robotic manipulation under limited observation(Fig.LABEL:fig:head_image).
+Our model employs a sequential dual-agent structure, with a Next-Best-View(NBV) agent to infer optimal viewpoints and a Next-Best-Pose(NBP) agent to determine gripper actions based on the observations from the NBV-inferred viewpoints(Fig.3).
+By alternating between sensor and motor actions, the model actively perceives and interacts with the environment to achieve task objectives.
+We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories.
+To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene grid to align with the camera viewpoint.
+We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions.
+We evaluate our model on 8 RLBench[9] tasks with limited viewpoints, demonstrating its effectiveness in viewpoint selection and action inference based on task requirements.
 The main contributions of this paper are as follows:
@@
 Active vision aims to enable agents to actively observe the environment based on specific task goals.
-It has been widely applied in tasks such as indoor visual navigation[12][15], object searching[16, 17, 18, 19] and 3D reconstruction[20].
-In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7][13][14][21][22] and for detecting fruit nodes in automated harvesting[23, 24, 25].
+It has been widely applied in tasks such as indoor visual navigation[12, 15], object searching[16, 17] and 3D reconstruction[18].
+In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7, 13, 14, 19] and for detecting fruit nodes in automated harvesting[20, 21, 22].
 Recent studies have extended the use of active vision to more general robotic manipulation tasks.
@@
 3D Voxel-based Manipulation.
-3D voxels are highly effective in representing spatial structures[26], making them widely used in domains such as medical imaging[27][28] and autonomous driving[29][30].
+3D voxels are highly effective in representing spatial structures[23], making them widely used in domains such as medical imaging[24, 25] and autonomous driving [26, 27].
 Recently, they have also been adopted in robotic manipulation models for visual representation.
@@
 Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
-Figure 2: Comparison of regular RL and asynchronous active vision-action decision processes.
-We introduce an intermediate time step $t^{\prime}$ between $t$ and $t+1$, dividing the interaction into NBV policy for viewpoint selection and the NBP policy for action selection.
-The task reward $r_{\text{t'}}$ for state $\mathbf{s}$ is derived from the reward $r_{\text{t+1}}$ after the NBP interaction, which enables sensor-motor joint training through shared task rewards.
+Figure 2: Decision process comparison.We divide single step interaction into NBV for viewpoint and NBP for action selection.
+Reward $r_{\text{t'}}$ is derived from $r_{\text{t+1}}$ after NBP interaction, enabling sensor-motor joint training through shared task rewards.
 Attention-driven Robotic Manipulation.
@@
 Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
-III background
 Figure 3: Overview of our model pipeline.
-The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optmal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal.
+The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optimal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal.
 The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
-Problem Formulation.
-We divide one asynchronous vision-action time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy from $t^{\prime}\sim t+1$.
-Therefore the asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},\mathcal{O^{\prime}},P^{\prime},R,\gamma>$ As shown in Fig.2, $\mathcal{O}$ and $\mathcal{O^{\prime}}$ represent the observation spaces of the agent at times $t$ and $t^{\prime}$, respectively.
-The NBV policy $\pi_{\text{v}}$ determines the camera viewpoint action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, and obtains a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ through transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$.
-The NBP policy $\pi_{\text{g}}$ then determines the gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$ based on the new observation $\mathbf{o}_{t^{\prime}}$.
-Through the transition probability $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$, scene observation $\mathbf{o}_{t+1}$ can be obtained by camera at time $t+1$, and the environment provides the task reward $r_{t+1}\in R$.
-Our goal is to jointly learn the policies $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{g}}$ that maximize the expected sensor-motor joint reward for the given manipulation task, which is $\pi^{*}_{\text{v}},\pi^{*}_{\text{g}}=\arg\underset{\pi_{\text{v}},\pi_{\text{p}}}{\max}\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
-Observation and Action Space We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\mathcal{O^{\prime}}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid voxelized from the scene RGB-D image captured by the camera, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state.
-Specifically, the raw RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, which is then voxelized through function $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center coordinates.
-The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and a occupancy flag for each voxel.
+III Problem Formulation
+We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$.
+The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},P^{\prime},r,\gamma>$, as shown in Fig.2.
+where $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{c}}},\mathcal{A^{\text{g}}}$ is the NBV and NBP action space, respectively.
+At time $t$, the NBV policy $\pi_{\text{v}}$ selects a camera action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ via transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$.
+The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$, with the next scene observation $\mathbf{o}_{t^{\prime}}$ obtained through $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$.
+The environment provides a reward $r_{t+1}\in\mathbb{R}$.
+Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
+Observation and Action Space.
+We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid derived from the RGB-D image, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state.
+Specifically, the RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, and voxelized via $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center.
+The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and an occupancy flag for each voxel.
 The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
-Following common settings[14][21][22], we restrict the camera viewpoint’s movement to the upper hemisphere at a fixed radius of $r^{\text{v}}=1.2\text{m}$ centered on the table, illustrated in Fig.8.
-The camera’s optical axis is assumed to always be directed towards the table center thus the viewpoint position can be described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$.
-We than define the NBV policy’s action as $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}\in\mathcal{A}^{\text{c}}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene.
-The NBP policy’s action is defined as $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}\in\mathcal{A}^{\text{g}}$, where $\mathbf{k}^{\text{g}}$ represents the discrete 6-DoF gripper pose index, and $d$ is the goal distance between the parallel gripper tips.
+Following common practice[14, 19], the camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, with the optical axis directed at the sphere center, illustrated in Fig.8(b).
+The viewpoint is described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$.
+The NBV policy action is $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene.
+The NBP policy action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}$ consists of the discrete 6-DoF gripper pose index $\mathbf{k}^{\text{g}}$ and the goal distance $d$ between the gripper tips.
 IV Method
-In this section, we provide a detailed explanation of the fundamental structure of our active vision-action model, along with three innovative strategies designed to enhance both model performance and learning efficiency.
 IV-A Model Structure
-Fig3 illustrates the basic pipline of our model.
-At time $t$, the voxelized image observation and proprioceptive state undergo a coordinate transformation for viewpoint alignment (Section IV-B), resulting in the processed observation $\mathbf{o}_{t}$.
-The NBV policy network, parameterized by $\Theta$, uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1).
-Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, representing the viewpoint that would best improve task success.
-The camera is then adjusted to the goal viewpoint $\mathbf{v}^{*}$, which is derived from the selected $\mathbf{k}^{\text{v}}$.
+We illustrates the basic pipline of our model in Fig3.
+Once the alignment process is completed(Section IV-B), the observation $\mathbf{o}_{t}$ is obtained.
+The DQN-based NBV policy network, parameterized by $\Theta$, then uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1).
+Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, which represents the viewpoint most likely to enhance task success.
+The camera is subsequently adjusted to the goal viewpoint $\mathbf{v}^{*}$, derived from the selected $\mathbf{k}^{\text{v}}$.
 $\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
-Based on the predicted ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the raw image observation from the viewpoint is cropped and aligned to generate the corresponding ROI 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$.
-This enables us to construct the observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$.
-Using the NBP policy network parameterized by $\Phi$, we infer the gripper closure action $d$ and the discrete gripper pose index $\mathbf{k}^{\text{g}}$ which is then transformed into the goal pose $\mathbf{g}^{*}$.
-Finally, a complete sensor-motor interaction step is achieved after gripper exectuation and rewarded by $r_{t+1}$.
+Based on the ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the image from $\mathbf{v}^{*}$ is cropped and aligned to generate the 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$ for the ROI.
+The NBP policy network, also DQN-based and parameterized by $\Phi$, infers the gripper closure $d$ and pose index $\mathbf{k}^{\text{g}}$ based on $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$.
+Finally, a complete sensor-motor interaction step is achieved after gripper interaction and rewarded by $r_{t+1}$.
 $\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
-We denote the sub-transition for the NBV policy as $\text{T}_{t}^{\text{v}}=<\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}}>$ and for the NBP policy as $\text{T}_{t}^{\text{p}}=<\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1}>$.
-Since the NBV and NBP policies share the same reward $r_{t+1}$, a complete interaction transition is denoted as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$.
-Upon completing an interaction episode, the collected transition is stored in the replay buffer together with the demonstration transitions.
-During model updates, both the NBV and NBP policies are updated using the same reward, thereby achieving sensor-motor joint learning.
-We employ a 3D-CNN based U-Net[35] as the image encoder for both policy networks, and three-layer fully connected network (FCN) for decoding the viewpoint and gripper actions.
+We denote a complete interaction transition as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$, where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}})$ represents the NBV policy’s sub-transition, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1})$ represents the NBP policy’s sub-transition.
+After each training episode, transitions are stored in the replay buffer along with demonstration data.
+Both policies are updated using the same task reward $r_{t+1}$, thereby achieving sensor-motor joint learning.
+Each policy employs a separate 3D-CNN based U-Net[32] as the image encoder, followed by a three-layer fully connected network(FCN) for action decoding.
 […]
 IV-B Viewpoint-centric Voxel Alignment
-Viewpoint movement expands the image observation space, increasing the variance of interaction samples[3].
-Specifically, images captured from a static viewpoint are primarily influenced by the scene state, following the mapping $f^{\text{static}}:\mathcal{S}\rightarrow\mathcal{O}^{\text{v}}$.
-In contrast, active viewpoint movement introduces an additional influence from the viewpoint itself, resulting in the mapping $f^{\text{active}}:\mathcal{S}\times\mathcal{V}\rightarrow\mathcal{O}^{\text{v}}$.
-Since the 3D-CNN lacks spatial rotation invariance, voxelizing the scene using the world coordinate system as the reference frame can cause features of similar structures from different viewpoints to appear distant in the feature space after encoding, as shown in Fig.4.
-This misalignment negatively impacts the overall training efficiency of the model.
 […]
@@
 IV-C Viewpoint-aware Demo Augmentation
+To provide early guidance and reduce exploration, we manually alternate camera and gripper movements to generate a small set of demonstration trajectories.
+Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of observation-action pairs collected at a fixed rate.
+To further leverage limited demo data, existing augmentation methods [1, 2, 28] first select key frame indices from the demo trajectory, construct transition samples from the key observation-action pairs, and then resample between key frames to generate additional transitions.
+However, they cannot handle the observation variance caused by viewpoint shifts.
+We define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, identifying key frames for both viewpoint and gripper movements.
+Gripper key frames $k_{j}^{\text{g}}$ are marked when the gripper is stationary or its closure state changes, while key viewpoints are defined as the first stable viewpoint between two gripper key frames, indexed by $k_{j}^{\text{c}}$.
+The key frame indices satisfy $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1.
+As shown in Fig.5, raw transition samples $\text{T}_{j}$ are then constructed as formalized in Eq.(5).
+$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{(\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}),(\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}),r_{j}\}\end{split}$ (5)
+To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$.
+The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
+$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ \end{split}$ (6)
+Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
 Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations(represented by hatched circles) between both viewpoint key frames(solid blue circles) and gripper key frames(solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
-To provide early guidance and reduce the need for extensive exploration, we manually alternate moving the camera and gripper to generate and collect a small number of demonstration trajectories as high-quality examples.
-Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of “observation-action” pairs collected at a fixed sampling rate.
-Existing demonstration augmentation methods[1][2][31] typically involve two main processes: key frame discovery and observation sampling.
-In the key frame discovery process, a limited set of key frame indices is selected from a trajectory $\tau_{d}$, and transition samples are constructed based on the corresponding “observation-action” pairs.
-The observation sampling process then generates additional transition samples by resampling the initial observations between the identified key frames.
-However, existing methods do not account for viewpoint movement.
-Our method incorporates key frame discovery and observation resampling, taking into account viewpoint movement, enabling viewpoint-aware demonstration augmentation.
-Specifically, we define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, which retrieves the key frames of both viewpoint and gripper movements in the demonstration trajectory based on predefined rules.
-The discovery of gripper key frames $k_{j}^{\text{g}}$ follows the method described in [1].
-We define key viewpoints as the first stable viewpoint between every two gripper key frames, corresponding to the frame index $k_{j}^{\text{c}}$.
-The relationship between key frame indices satisfies $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1.
-As shown in Fig.5, after the key frame discovery process, we construct the raw transition samples $\text{T}_{j}$ formalized in Eq.(5).
-$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{<\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}>,<\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (5)
-To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented frame indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$.
-The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
-$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ &=\{<\mathbf{o}_{(k_{j-1}^{\text{g}})^{\text{aug}}},\mathbf{a}_{(k_{j-1}^{\text{g}})^{\text{aug}}}^{\text{c}},\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}}>,\\ &<\mathbf{o}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{a}^{\text{g}}_{(k_{j}^{\text{c}})^{\text{aug}}},\mathbf{o}_{k_{j}^{\text{g}}}>,r_{j}\}\end{split}$ (6)
-Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
-Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene.
-We encourage the model to select non-empty ROIs (green dashed boxes) and maximize the number of occluded voxels within the ROI that can be observed from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$ compared to the original viewpoint $\mathbf{v}_{t}$.
-This strategy promotes active interaction with the environment and enables the model to capture more structural information within the ROI.
 IV-D Task-agnostic Auxiliary Reward
-We use the task completion signal as a sparse task reward to avoid designing targeted rewards for different tasks; however, sparse reward only provides positive feedback to the model at the end of an episode, making it difficult to efficiently guide the model in sensor-motor joint learning.
-To address this, we introduce two task-agnostic auxiliary rewards that provide real-time feedbacks to the training process.
-To encourage effective interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, reward $r_{\text{i}}$ provides a positive signal when the NBV policy infers an ROI in a reachable and non-empty area of the scene, and penalizes actions that select an ROI in an empty or unreachable area due to distance or trajectory planning failures.
+We use sparse reward to avoid designing task-specific rewards.
+However, since sparse reward only provides feedback at the end of an episode, it hinders efficient sensor-motor learning.
+Here we introduce two task-agnostic auxiliary rewards to provide real-time guidance to the training process.
+To encourage interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, $r_{\text{i}}$ provides positive signals when the NBV policy selects a reachable, non-empty ROI and penalizes actions targeting empty or unreachable areas due to execution failures.
 The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
-$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI not reachable and empty}\end{cases}$ (7)
-To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward , denoted as $r_{\text{e}}$, inspired by [24].
-The core idea is to reduce the occupancy entropy within the ROI after adjusting the viewpoint, as illustrated in Fig.6.
-Specifically, assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel is calculated using the Shannon entropy function, as shown in Eq.(8), where $p_{\text{o}}(i)$ represents the voxel’s occupancy probability.
-If a voxel is occluded, its occupancy probability is set to $p_{\text{o}}=0.5$.
+$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}\end{cases}$ (7)
+To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward($r_{\text{e}}$), inspired by [21].
+The goal is to reduce occupancy entropy within the ROI after viewpoint adjustment, as shown in Fig.6.
+Assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel($e_{i}$) is calculated using the Shannon entropy function(Eq.(8)), where $p_{\text{o}}(i)$ represents its occupancy probability.
+If a voxel is occluded, $p_{\text{o}}$ is set to 0.5.
 $e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
@@
 $\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
-We define the reward $r_{\text{e}}$ as Eq.(10), where the coefficient $k$ is empirically set to 1.0.
-$r_{\text{e}}=k\cdot\Delta E,k=1.0$ (10)
-The auxiliary reward $r_{\text{i}}$ necessitates the cooperation of both the NBV and NBP policies, whereas $r_{\text{e}}$ is exclusive to the NBV policy.
-Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, while the reward for updating the NBV agent is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion.
-It is important to note that the introduced rewards are designed to be task-agnostic, which ensures that the rewards can generalize across different tasks without the need for manual tuning.
-Figure 7: Reletive task occlusion rate of selected tasks.
-We use the active viewpoint occlusion rate of 30 demonstration trajectories for each task as a benchmark.
-Based on this, we calculate the relative task occlusion rates for the front, overhead, and wrist viewpoints in the demonstration trajectories of each task.
+We define reward $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0.
+The auxiliary reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy.
+Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion.
+These task-agnostic rewards generalize across tasks without the need for manual tuning.
+Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene.
+The model is encouraged to select non-empty ROIs and reduce occluded voxels from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$.
 V Experiment Setup
-V-A Task Setting
-We trained and evaluated our model on RLBench [9], a public benchmark designed for reinforcement learning manipulation algorithms.
-When selecting evaluation tasks, we considered two primary criteria: 1) Single task: each task must have at least one fixed viewpoint with a high level of visual occlusion.
-2) Task range: no single fixed or wrist-mounted camera can provide a low occlusion rate for all tasks.
-Since the region of interest (ROI) the robot focuses on may change during different manipulation phases, and there may be no clear manipulation objects (e.g., when the gripper is transferring in midair), it is challenging to use an image segmentor to locate objects and define the task occlusion rate as in previous methods[14][24].
-To address this, we define the Task Occlusion Rate (TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(11), Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation.
-To compare the occlusion levels across different tasks, we assume that the collected demo trajectories contain the strongest task-specific prior knowledge.
+Task Setting.
+We trained and evaluated our model on RLBench[9], a widely-used public benchmark for robotic manipulation in reinforcement learning.
+Task selection followed two criteria: 1) each task must include at least one viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across selected tasks.
+As the ROI shifts throughout the manipulation process, and there may be phases without distinct objects(e.g., midair gripper movements), defining occlusion rates based on object segmentation, as done in previous methods[14, 21], is impractical.
+To address this, we define Task Occlusion Rate(TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(10).
+Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation.
+To compare the occlusion levels across selected tasks, we assume that the human-demoed active viewpoint contains the strongest task-specific prior knowledge.
 Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint(oracle).
 $\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (10)
-Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under common overhead, front , wrist-mounted cameras as well as the human-demoed active camera, based on $100$ demo trajectories.
-The results show that, due to task characteristics and initial scene randomization, both fixed and wrist viewpoints experience varying degrees of occlusion across the 8 tasks.
-In contrast, the active camera demonstrates the potential to achieve lower task occlusion.
-V-B Baselines
-We compaired our model against two categories baseline models: 1) visual manipulaiton model that utilze common used static third-person cameras and first-person wrist-mounted camera; 2) Active third-person camera based manipulation models.
-For the first category, we compared our active vision-action model against the recent visual-based C2FARM [2] model with static front, overhead camera as well as the wrist-mounted camera settings.
-This comparison aims to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
-For the second category, using the same third-person active camera setting, we firstly compared our model against the Behavior Cloning (BC) model.
-Through this comparison, we seek to investigate whether our model not only mimics the given limited demonstrations but also learns new sensor-motor skills through interaction with the environment.
-Additionally, we tested a Random viewpoint model, following [3], which shares the same architecture as ours but with the learning of the NBV policy network disabled (learning rate set to 0).
-This setup allows us to evaluate the importance of the active viewpoint selection in the NBV policy and to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random viewpoint choices.
-V-C Training Setting
-We first collected 30 demo trajectories for each task and stored them in the replay buffer as prior samples or the training dataset, and all models were trained from scratch.
-The implementation of C2FARM was based on their official release and trained using sparse task rewards.
-The Behavior Cloning model was trained with the Mean Squared Error (MSE) loss, while the Random model utilized the same reward structure as our model, as described in SectionIV-D.
-All models were trained using the Adam optimizer with an initial learning rate of $0.0005$, with no learning rate decay applied during training.
-Each model was trained for 40,000 updates with a batch size of 128, and all experiments were conducted on a machine equipped with 4 NVIDIA GTX 4090 GPUs.
-TABLE I: Task success rate: Our model outperform baselines on most tasks and on average . Mean and std of 100 tests are reported.
+Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under overhead, front, wrist-mounted cameras as well as the human-demoed active camera.
+Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 100 demonstration trajectories for each task as a benchmark.
+Baselines.
+We compared our model against two baseline categories: 1) visual manipulation model that utilize static third-person and wrist-mounted cameras, and 2) active third-person camera-based models.
+For the first category, we compared our model against C2FARM [2] with static front, overhead and wrist-mounted camera settings to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
+For the second category, we compared our model with Behavior Cloning(BC) to test whether it can learn new sensor-motor skills beyond demonstration imitation.
+We also tested a Random viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled(learning rate set to zero).
+This setup allows us to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random choices.
+Training Setting.
+We collected 30 demo trajectories for each task and used them as prior knowledge for training.
+C2FARM was implemented based on its official release and trained with sparse task rewards.
+The Behavior Cloning model was trained with MSE loss, and the Random model used the same reward structure as our model(SectionIV-D).
+All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates, with a batch size of 128.
+Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
+TABLE I: Task success $\uparrow$ . Mean and std of 100 tests are reported(3 seeds).
 Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
@@
 Ours(active) 0.81±0.02 0.56±0.04 0.73±0.04 0.50±0.03 0.85±0.03 0.96±0.01 0.78±0.02 0.92±0.01 0.76±0.15
-TABLE II: Task episode length. The maximum length of each task episode is limited to 10 steps. Mean and std of 100 tests are reported.
+TABLE II: Task episode $\downarrow$ . The maximum length of each task episode is limited to 10 steps.
 Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
@@
 Main Performance.
-We first analyzed the average task success rate and the average episode length across different algorithms, which provide a macro-level assessment of the models’ proficiency in learning sensor-motor coordination.
-Table I presents the average success rate for tasks across 100 testing episodes under 3 evaluation seeds.
-Our model achieved the highest average success rate across the eight tasks, outperforming the second-best by 21%.
-It also exhibited relatively high manipulation performance in scenarios with significant occlusion from fixed viewpoints, such as meat_on_grill, meat_off_grill and unplug_charger.
-Similarly, Table II shows that the average episode length for tasks follows the same trend as the success rate.
-Our model achieved the shortest average manipulation steps across all tasks, outperforming the second-best model by 28.2%.
-Furthermore, it demonstrated the lowest variance in manipulation steps, highlighting the model’s high proficiency and consistency after training.
-Moreover, we observed that in tasks where the occlusion rate is low or the scene object positions are relatively predictable, fixed and wrist-mounted cameras benefit from smaller variance in observation samples, leading to enhanced skill acquisition.
-For example, in tasks such as basketball_in_hoop, press_switch, put_rubbish_in_bin, the C2FARM model achieved comparable success rates from both front and overhead views, occasionally even surpassing our model.
-However, it is important to note that when these fixed cameras fail to capture adequate observations, performance tends to decline sharply, leading to higher performance variance across the 8 tasks.
-In contrast, our model is less sensitive to visual occlusion and consistently maintains strong manipulation performance.
-Under the active viewpoint setting, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), its effectiveness was constrained by the limited number of training samples, which restricted its capacity to learn interaction skills.
-As a result, it frequently failed to complete the assigned tasks, achieving an average success rate of only 14%.
-Meanwhile, the random viewpoint model performed poorly across all 8 tasks, exhibiting the weakest manipulation performance, indicating its inability to effectively learn sensor-motor coordination skills.
-Figure 8: Viewpoint distribution overview: Red circles indicate the distribution of manually set viewpoints across 30 demo trajectories, while blue triangles represent the distribution of viewpoint positions during successful episodes of our model over 100 test episodes.
-The model’s inferred viewpoints do not strictly mimic the demonstration trajectories but instead exhibit a degree of refinement and generalization.
-Figure 9: Partial rollout cases of our model.
-The figure presents the image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the associated gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
-TABLE III: A comparison of information gain to baselines.
-      Models              ASSIG       (avg/max)              AIIG       (avg/max)
-      C2FARM(wrist)       0.0/0.0       N/A
-      C2FARM(overhead)       0.0/0.0       0.0/0.0
-      C2FARM(front)       0.0/0.0       0.0/0.0
-      Random(active)       0.0/5.1%       0.0/5.2%
-      BC(active)       0.7%/2.3%       5.6%/18.1%
-      Ours(active)       3.5%/6.9%       6.8%/25.1%
+We first analyzed the average task success rate(SR) and episode length(EL) across different models to evaluate overall sensor-motor performance.
+Table I shows that our model achieved the highest success rate, outperforming the second-best by 21%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger.
+Similarly, Table II shows that our model had the shortest average manipulation steps, outperforming the second-best by 28.2% and showing the least variance in step counts.
+Fixed and wrist-mounted cameras performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, C2FARM occasionally matching or surpassing our model.
+However, their performance dropped sharply in occluded scenarios, leading to a higher variance.
+In contrast, our model remained robust in occluded tasks.
+Although the BC model demonstrated some ability to adjust its viewpoint(as detailed below), it struggled with limited training samples, achieving only a 14% success rate.
+The random viewpoint model performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
+Figure 8: Partial rollout cases and viewpoint distribution.
+(a) The image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
+(b) Red circles indicate manually set viewpoints from 30 demo trajectories, while blue triangles represent the learned viewpoint positions from successful episodes across 100 test runs.
 Active Perception Strategy Analysis.
-We further analyzed the perception strategies of all models during the operation process using two metrics: Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG).
-These metrics measure the models’ ability to acquire observation information through multi-step and single-step interactions, respectively.
-Specifically, we define ASSIG as the average reduction in entropy within the ROI before and after a single viewpoint movement.Similarly, the initial information gain is defined as the average reduction in entropy of the ROI after a viewpoint movement, compared to the initial viewpoint.
-These two metrics are given by:
+We further analyzed the perception strategies of all models using two metrics: Average Single Step Information Gain(ASSIG) and Average Initial Information Gain(AIIG).
+ASSIG captures the average reduction in ROI entropy before and after a single viewpoint shift, while AIIG measures the entropy reduction compared to the initial viewpoint.
 $\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (11)
 $\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (12)
-Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the number of steps required to complete episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9).
+Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the length of episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9).
 The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint.
-The statistical results in Table III show that our model significantly improves the observation of the ROI region by independently adjusting camera viewpoints, outperforming the second-best BC model by 2.8% and 1.2% in the ASSIG and AIIG scores, respectively.
-In contrast, static front and overhead cameras, which are unable to move, fail to achieve either single-step or initial information gain through viewpoint shifts.
-Similarly, the wrist-mounted camera, being unable to move independently of the gripper, is limited to inferring the next gripper actions based on the current viewpoint.
-This results in a zero ASSIG score and makes it unsuitable for AIIG computation.
-Fig.8 illustrates the distribution of viewpoints learned by our model alongside the key viewpoints distribution from 30 demo trajectories.
-Notably, the viewpoints learned by our model do not strictly follow the patterns observed in the demonstration trajectories.
-For tasks such as put_rubbish_in_bin, press_switch, unplug_charger, and meat_off_grill, the learned viewpoints are more concentrated.
-In contrast, for tasks like meat_on_grill, open_drawer, and take_money_out_safe, the model explores a broader range of viewpoints.
-This variation indicates that our model has effectively learned to select task-relevant viewpoints through interactive learning, rather than merely replicating the active perception patterns from the demonstration data.
-Furthermore, Fig.
-8 presents camera observations at keyframe moments of partial task episodes, showcasing our model’s capacity for active visual perception during task execution.
-Figure 10: Comparison of success, timeout, and execution failure rate across models.
+As shown in Table III, our model significantly enhances ROI observation by adjusting viewpoints, outperforming the BC model by 2.8% in ASSIG and 1.2% in AIIG.
+In contrast, static front and overhead cameras, which cannot shift viewpoints, fail to achieve information gain.
+The wrist-mounted camera, unable to move independently of the gripper, shows zero ASSIG and is unsuitable for AIIG computation.
+Fig.8(b) compares the distribution of viewpoints learned by our model with the manually set active viewpoints from the demo trajectories used during training.
+Unlike demo patterns, our model’s learned viewpoints are more concentrated in some tasks(Basketball_in_hoop, unplug_charger) and more varied in others(meat_on_grill, open_drawer), indicating it learned to select task-relevant viewpoints rather than simply imitating demo data.
+The camera observations at key moments(Fig.
+8) further showcase our model’s active visual perception capabilities.
+Figure 9: Proportion of success, timeouts, and execution failures across test episodes for different models.
+TABLE III: Comparison of information gain and interaction capacity across different models.
+Models ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
+C2FARM(wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
+C2FARM(overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
+C2FARM(front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
+Random(active) 0.0/5.1% 0.0/5.2% 9.07±1.47 9.49±1.07
+BC(active) 0.7%/2.3% 5.6%/18.1% 6.31±2.46 7.54±2.15
+Ours(active) 3.5%/6.9% 6.8%/25.1% 1.64±0.61 2.15±1.03
 Environmental Modeling and Interactive Capabilities.
-To further evaluate the environment modeling and interaction capabilities of various models, we recorded the success rate, timeout rate, and execution failure rate over 100 test episodes for each model, as illustrated in Fig.10.
-The C2FARM model showed a high likelihood (up to 42%) of episodes terminating prematurely due to failures in low-level trajectory execution.
+We evaluated the models’ environment modeling and interaction capabilities by recording success, timeout(TO), and execution failure(EF) rates over 100 test episodes(Fig.9).
+The C2FARM model exhibited a high failure rate(up to 42%) due to trajectory execution failures.
 Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors.
-While execution failures accounted for only 4% of the BC model’s episodes, as many as 82% of the episodes failed due to timeouts.
-However, our model showed only a 13% rate of episodes ending prematurely due to trajectory execution failures and a mere 9% failure rate due to timeouts.
-These results suggest that our model effectively integrates low-level motion execution considerations when inferring manipulation actions, resulting in more robust performance across the test episodes.
-Additionally, as shown in Table IV, we analyzed the average steps to first interaction with scene objects and the average number of non-interactive steps during testing for each model.
-Our findings indicate that the behavior cloning exhibited a relatively high number of non-interactive steps, which aligns with the high timeout rate shown in Fig.10.
-This pattern suggests that behavior cloning struggles to learn precise interactive manipulations from a limited number of demonstration trajectories, leading to frequent non-interactive steps and episode timeouts.
-In contrast, our model consistently initiates interaction with the environment more quickly and significantly reduces the time spent on non-interactive actions.
-These results demonstrate that, after training, our model achieves superior accuracy in modeling environmental dynamics and understanding structural relationships.
-TABLE IV: First interaction step and non-interaction step comparison.
-Models First Interaction Step Non-interaction Step
-C2FARM(wrist) 3.21±1.82 4.86±2.70
-C2FARM(overhead) 2.97±1.70 4.41±2.77
-C2FARM(front) 3.41±2.28 4.17±2.74
-Random(active) 9.07±1.47 9.49±1.07
-BC(active) 6.31±2.46 7.54±2.15
-Ours(active) 1.64±0.61 2.15±1.03
-Ablation Study In this section, we investigate the impact of three key strategies on manipulation performance and various evaluation metrics: (1) viewpoint-centric voxel alignment (Section IV-B), (2) viewpoint-aware demon augmentation (Section IV-C), and (3) task-agnostic auxiliary rewards (Section IV-D).
-To assess the individual and combined contributions of these strategies, we conducted separate ablation tests for each, followed by a comprehensive test removing all three strategies.
-As shown in Table V, each of the three strategies contributed to notable improvements in the success rate of our proposed model.
-Among the three, voxel alignment and viewpoint-aware demonstration augmentation had the most substantial impact, with the success rate dropping by approximately 40% when either of these strategies was ablated, highlighting their critical roles.
-The auxiliary reward strategy, while less significant than the other two, still provided approximately a 6% improvement, demonstrating its contribution to overall performance.
-Beyond success rate improvements, all three strategies contributed to reductions in episode length, overtime ratio, planning failure rate, and the average number of non-interactive steps per episode.
-This demonstrates their broader importance to the algorithm’s effectiveness.
-Interestingly, while the voxel alignment strategy significantly boosts overall performance, it leads to a reduction in ASSIG score.
+Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts.
+In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
+As shown in Table III, we also analyzed the average steps to first interaction(FI) and non-interactive steps(NI) per model.
+The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise manipulations from limited demonstrations.
+In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
+Ablation Study.
+We evaluated the impact of three strategies on manipulation performance:(1) viewpoint-centric voxel alignment(Section IV-B), (2) viewpoint-aware demo augmentation(Section IV-C), and (3) task-agnostic auxiliary rewards(Section IV-D).
+Separate ablation tests were conducted for each, as well as a combined test removing all strategies.
+Table IV shows that each strategy effectively improved success rates(SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
+Beyond success rate improvements, all three strategies contributed to reductions in episode length(EL), timeouts rate(TO), execution failures(EF), and the average number of non-interactive steps(NI) per episode.
+This demonstrates their broader importance to the model’s effectiveness.
+Interestingly, while the voxel alignment strategy significantly boosts overall performance, we found it leads to a reduction in ASSIG score.
 We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations.
 However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
-When all three strategies were ablated, the performance across the eight tasks showed the most severe decline, with the average success rate dropping to just 22%, even lower than the static-viewpoint baseline.
-This highlights the synergistic benefits of these three strategies, as their combination significantly boosts the model’s overall performance.
-TABLE V: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards
-Models Success Rate Episode Length Overtime Rate Planning Failed Rate
-Ours(w/o align) 0.34±0.41 7.61±2.94 0.14±0.12 0.53±0.39
-Ours(w/o aug) 0.31±0.37 7.99±2.44 0.15±0.17 0.55±0.37
-Ours(w/o aux) 0.70±0.19 5.52±1.51 0.15±0.10 0.16±0.12
-Ours (w/o align+aug+aux) 0.22±0.34 8.46±2.41 0.22±0.22 0.57±0.34
-Ours 0.76±0.15 4.88±1.16 0.10±0.06 0.15±0.11
-Models First Interaction Non-interaction AIIG ASSIG
-Ours(w/o align) 4.33±2.80 5.81±3.30 0.046±0.05 0.082±0.07
-Ours(w/o aug) 4.49±3.34 5.74±3.58 0.014±0.05 0.020±0.03
-Ours(w/o aux) 1.85±0.96 2.38±1.25 0.067±0.08 0.028±0.02
-Ours (w/o align+aug+aux) 5.74±3.37 6.72±3.28 0.036±0.07 0.083±0.06
-Ours 1.64±0.61 2.15±1.03 0.068±0.09 0.035±0.03
-VII Conclusion and Feature Work
+Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.
+TABLE IV: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards.
+Ablation Setting Result Proportion Interactivity Sensing
+align aug aux SR EF TO EL FI NI AIIG ASSIG
+0.34 0.53 0.14 7.61 4.33 5.81 0.046 0.082
+0.31 0.55 0.15 7.99 4.49 5.74 0.014 0.020
+0.70 0.16 0.15 5.52 1.85 2.38 0.067 0.028
+0.22 0.57 0.22 8.46 5.74 6.72 0.036 0.083
+0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035
+VII Conclusion and Future Work
 In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model.
-Our model first infers the optimal observation viewpoint accroding to the task goal and then determines the corresponding manipulation action based on the updated viewpoint’s obsevation.
-To enhance performance and stability, our model integrates three strategies: voxel alignment for spatial consistency, demonstration augmentation for better learning, and auxiliary rewards for real-time feedback.
-Together, these strategies help the model handle occlusions and improve sensor-motor coordination.
-Experimental results show that, compared to passive observation-based model, our model significantly improves manipulation performance through actively observation.
-It further learns sensor-motor coordination through interaction with the environment, rather than being restricted to a limited set of demonstration trajectories.
-Despite these promising results, the model’s performance can still be affected by complex occlusions and dynamic environments.
-Future research could explore incorporating real-time visual feedback to better handle dynamic changes in the environment.
+Our model first infers the optimal viewpoint based on the task goal, then selects the manipulation action from the updated observation.
+To enhance performance and training stability, we integrate voxel alignment for spatial consistency, demonstration augmentation for enhanced learning, and auxiliary rewards for real-time training guidance.
+These strategies enable the model to handle occlusions and improve sensor-motor coordination.
+Experimental results show that our model significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment.
+However, complex occlusions and dynamic environments remain challenging.
+Future work could explore real-time visual feedback to better adapt to environmental changes.

### v3

@@
 Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
+Index Terms:
+Manipulation Planning; Reinforcement Learning; Perception for Grasping and Manipulation.
 I Introduction
 Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3].
-While these cameras reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
-However, humans can adjust their head and eye positions based on task goals.
-This helps them obtain advantageous visual information for action decision [5].
-Inspired by this, recent active vision-action models enable agents to actively adjust their sensory viewpoints while interacting with the environment[3, 6, 7, 8], mimicking human capabilities.
-Active vision-action models, like human perception, can be categorized as synchronous or asynchronous[10].
+While these setups reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
+Alternatively, humans can adjust their head and eye positions according to task goals, which helps them obtain advantageous visual information for action decision-making [5].
+Inspired by this, recent active vision-action models allow agents to actively adjust their sensory viewpoints while interacting with the environment [3, 6, 7, 8].
+These models, which mimic human capabilities, can be categorized as either synchronous or asynchronous [9].
 Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 10, 11].
@@
 We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories.
-To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene grid to align with the camera viewpoint.
+To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene voxel grid to align with the adjusted viewpoint.
 We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions.
@@
   • We introduce methodological innovations, including viewpoint-aware demo augmentation, viewpoint-centric voxel alignment, and a task-agnostic auxiliary reward strategy, to enhance model performance, learning efficiency, and training stability.
-  • Comprehensive experiments on public RLBench tasks validate the effectiveness and robustness of our approach.
+  • Comprehensive experiments on 8 public RLBench tasks validate the effectiveness and robustness of our approach.
 II Related Work
 Active Vision for Manipulation.
-Active vision aims to enable agents to actively observe the environment based on specific task goals.
-It has been widely applied in tasks such as indoor visual navigation[12, 15], object searching[16, 17] and 3D reconstruction[18].
-In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7, 13, 14, 19] and for detecting fruit nodes in automated harvesting[20, 21, 22].
-Recent studies have extended the use of active vision to more general robotic manipulation tasks.
-For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes.
-Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects.
-Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints.
-Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills.
-Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
+Active vision [15] aims to enable agents to actively observe their environment based on specific task goals, such as visual navigation[11], object recognition[16, 17], and 3D reconstruction [18, 19].
+In robotic manipulation, active vision is primarily applied to tasks such as object grasping [7, 13, 14, 20] and fruit harvesting [21, 22].
+For more general manipulation tasks, Cheng et al.
+[10] trained an RL-based viewpoint policy to detect occluded objects and perform a pushing task.
+Lv et al.
+[8] trained a model-based active sensing-action model for flexible object manipulation.
+Chuang et al.
+[23] used imitation learning to train an active viewpoint strategy that provides low-occlusion observations for peg insertion tasks.
+Shang et al.
+[6] developed a self-supervised reward mechanism to jointly learn active visual sensing and motor skills for door opening and table wiping tasks.
+Dengler et al.
+[24] trained an active viewpoint policy to learn how to push objects in confined spaces.
+Our work is most similar to [10], but unlike their model, we do not rely on a task-specific object detector.
+Furthermore, our active vision-action model uses an RL-based dual-agent asynchronous decision-making structure, allowing the agent to observe the environment first and then interact with it, similar to human behavior.
 3D Voxel-based Manipulation.
@@
 Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
-Figure 2: Decision process comparison.We divide single step interaction into NBV for viewpoint and NBP for action selection.
+Figure 2: Decision process comparison.
+We divide single step interaction into NBV for viewpoint and NBP for action selection.
 Reward $r_{\text{t'}}$ is derived from $r_{\text{t+1}}$ after NBP interaction, enabling sensor-motor joint training through shared task rewards.
 Attention-driven Robotic Manipulation.
-The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs, improving both training efficiency and model performance[30].
-ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy 2D regions via DQN, while C2FARM predicts regions of interest(ROIs) in 3D space.
+The attention mechanism, widely applied in fields such as object detection and natural language processing, contributes to robotic manipulation by focusing on the most task-relevant scene information from raw image inputs[32], improving both training efficiency and model performance.
+Gualtieri et al.
+[33] generate grasp actions by applying attention to point cloud images to select noteworthy local regions and performing fine perception on these regions.
+ARM [1] and C2FARM [2] use attention mechanisms to focus on key visual input regions, with ARM selecting noteworthy regions from 2D inputs, while C2FARM predicts 3D regions of interest (ROI) from voxelized 3D images.
 Liang et al.
@@
 However, these methods largely rely on single or multiple fixed cameras from different viewpoints to capture images of the scene, lacking the capability to actively choose their viewpoints.
-Our model partially incorporates the attention mechanism from C2FARM and endows the model with the ability to predict the optimal task-related observation viewpoint, further mimicking the human’s active visual perception capability.
-Figure 3: Overview of our model pipeline.
-The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optimal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal.
-The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
+Our model incorporates the attention mechanism from C2FARM to select noteworthy local 3D regions from the scene but further enhances our model with the ability to predict the optimal observation angle for these regions, further mimicking human active visual perception capabilities.
+Figure 3: Illustration of model pipeline and action spaces.
+The NBV policy infers the 3D ROI position $\mathbf{r}^{*}$ based on the global scene observation and predict the optimal viewpoint $\mathbf{v}^{*}$ for the ROI observation according to the given task goal.
+The NBP agent determines the gripper actions based on the ROI observation from $\mathbf{v}^{*}$.
 III Problem Formulation
 We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$.
-The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},P^{\prime},r,\gamma>$, as shown in Fig.2.
-where $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{c}}},\mathcal{A^{\text{g}}}$ is the NBV and NBP action space, respectively.
-At time $t$, the NBV policy $\pi_{\text{v}}$ selects a camera action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ via transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$.
-The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$, with the next scene observation $\mathbf{o}_{t^{\prime}}$ obtained through $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$.
-The environment provides a reward $r_{t+1}\in\mathbb{R}$.
-Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
+The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{S},\mathcal{O},\mathcal{A^{\text{v}}},P,\mathcal{A^{\text{p}}},P^{\prime},r,\gamma>$, as shown in Fig.2, where $\mathcal{S}$ is the environment states, $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{v}}},\mathcal{A^{\text{p}}}$ is the NBV and NBP action space, respectively.
+At time $t$, the NBV policy $\pi_{\text{v}}$ selects a viewpoint action $\mathbf{a}_{t}^{\text{v}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{v}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\sim P(\cdot|\mathbf{s}_{t},\mathbf{a}_{t}^{\text{v}})$.
+The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{p}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{p}}}$, with the next scene observation $\mathbf{o}_{t+1}\sim P^{\prime}(\cdot|\mathbf{s}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{p}})$ and receives a reward $r_{t+1}\in\mathbb{R}$.
+Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\cdot r(\mathbf{s}_{t},\mathbf{a}^{\text{v}}_{t},\mathbf{a}^{\text{p}}_{t^{\prime}},\mathbf{s}_{t+1})\Big{]}$.
 Observation and Action Space.
-We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid derived from the RGB-D image, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state.
-Specifically, the RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, and voxelized via $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center.
-The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and an occupancy flag for each voxel.
-The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
-Following common practice[14, 19], the camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, with the optical axis directed at the sphere center, illustrated in Fig.8(b).
-The viewpoint is described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$.
-The NBV policy action is $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene.
-The NBP policy action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}$ consists of the discrete 6-DoF gripper pose index $\mathbf{k}^{\text{g}}$ and the goal distance $d$ between the gripper tips.
+The observation spaces $\mathcal{O}=\{\mathcal{O}^{\text{i}},\mathcal{O}^{\text{s}}\}$, where $\mathbf{o}^{\text{i}}\in\mathcal{O}^{\text{i}}$ represents the 3D voxelized scene image from the RGB-D data, and $\mathbf{o}^{\text{s}}\in\mathcal{O}^{\text{s}}$ is the proprioceptive state.
+Specifically, the 3D voxel grid $\mathbf{o}^{\text{i}}\in\mathbb{R}^{n^{2}\times(3+M+1)}$ contains the voxel coordinates, $M$-dimensional features, and one occupancy flag for each voxel, where $n=16$ and $M=11$.
+The proprioceptive state $\mathbf{o}^{\text{s}}=\{\mathbf{v},\mathbf{g},d\}$ comprises the viewpoint spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$, the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$ and the gripper finger tips distance $d\in\mathbb{R}$.
+The camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, illustrated in Fig.3, with the optical axis directed at the sphere center.
+The NBV action $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{Z}^{2}$ is the discrete viewpoint index for $\theta^{\text{v}}\in[10^{\circ},60^{\circ}]$ and $\phi^{\text{v}}\in[-135^{\circ},135^{\circ}]$ with a resolution of $5^{\circ}$.
+$\mathbf{k}^{\text{r}}\in\mathbb{Z}^{3}$ denotes the position index of the 3D ROI that requires attention within the workspace.
+The $1\text{m}^{3}$ workspace is discretized into $16^{3}$ uniform grids, resulting in each dimension of $\mathbf{k}^{\text{r}}$ having an index range of $[0,15]$.
+The NBP action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}}\}$ consists of the discrete gripper pose index $\mathbf{k}^{\text{g}}\in\mathbb{Z}^{6}$ and desired gripper closure state $\mathbf{k}^{\text{c}}\in[0,1]$.
+The gripper position is constrained within the $0.3\text{m}^{3}$ ROI, which is further discretized into $16^{3}$ grids.
+The gripper’s rotation is represented by Euler angles on each axis, discretized in the range $[0^{\circ},360^{\circ}]$ with a resolution of $5^{\circ}$.
+It is important to note that while both NBV and NBP agents share the same observation format, NBV observes the entire scene, whereas NBP focuses only on the 3D ROI image, achieving an attention effect [2].
 IV Method
 IV-A Model Structure
-We illustrates the basic pipline of our model in Fig3.
-Once the alignment process is completed(Section IV-B), the observation $\mathbf{o}_{t}$ is obtained.
-The DQN-based NBV policy network, parameterized by $\Theta$, then uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1).
-Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, which represents the viewpoint most likely to enhance task success.
-The camera is subsequently adjusted to the goal viewpoint $\mathbf{v}^{*}$, derived from the selected $\mathbf{k}^{\text{v}}$.
-$\mathbf{a}_{t}^{\text{c}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}})$ (1)
-Based on the ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the image from $\mathbf{v}^{*}$ is cropped and aligned to generate the 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$ for the ROI.
-The NBP policy network, also DQN-based and parameterized by $\Phi$, infers the gripper closure $d$ and pose index $\mathbf{k}^{\text{g}}$ based on $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$.
-Finally, a complete sensor-motor interaction step is achieved after gripper interaction and rewarded by $r_{t+1}$.
-$\mathbf{a}_{t^{\prime}}^{\text{g}}=\arg\underset{\mathbf{k}^{\text{g}},d}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},d)$ (2)
-We denote a complete interaction transition as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$, where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}})$ represents the NBV policy’s sub-transition, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1})$ represents the NBP policy’s sub-transition.
-After each training episode, transitions are stored in the replay buffer along with demonstration data.
-Both policies are updated using the same task reward $r_{t+1}$, thereby achieving sensor-motor joint learning.
-Each policy employs a separate 3D-CNN based U-Net[32] as the image encoder, followed by a three-layer fully connected network(FCN) for action decoding.
-[…]
+Both the NBV and NBP policies are implemented using a DQN network with a 3D U-Net backbone[35], parameterized as $\Theta$ and $\Phi$, respectively.
+The NBV agent infers the center position index $\mathbf{k}^{\text{r}}$ of the ROI in the scene and the optimal viewpoint index $\mathbf{k}^{\text{v}}$ for the ROI region based on the preprocessed aligned observation $\mathbf{o}_{t}$, as shown in Eq.(1).
+It then interacts with the environment and moves to move the viewpoint to the corresponding goal position $\mathbf{v}^{*}$.
+$\mathbf{a}_{t}^{\text{v}}=\arg\underset{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}}}{\max}Q_{\Theta}(\mathbf{o}_{t},\mathbf{k}^{\text{v}},\mathbf{k}^{\text{r}})$ (1)
+The depth image data captured at time $t^{\prime}$ is cropped and voxelized, yielding the voxel grid observation $\mathbf{o}^{\text{i}}_{t^{\prime}}$ for the ROI region under the new viewpoint.
+The NBP agent then infers the gripper pose index $\mathbf{k}^{\text{g}}$ and closure action $\mathbf{k}^{\text{c}}$ based on observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{i}}_{t^{\prime}},\mathbf{o}^{\text{s}}_{t^{\prime}}\}$.
+$\mathbf{a}_{t^{\prime}}^{\text{p}}=\arg\underset{\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}}}{\max}Q_{\Phi}(\mathbf{o}_{t^{\prime}},\mathbf{k}^{\text{g}},\mathbf{k}^{\text{c}})$ (2)
+Finally, the arm trajectory is planned using an off-the-shelf trajectory planner and executed through multiple control steps to move the gripper to the goal position $\mathbf{g}^{*}$, completing one sensor-motor interaction and receiving the reward $r_{t+1}$.
+The transition of the global agent denote as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$ , where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{v}},\mathbf{o}_{t^{\prime}})$ is the local transition of the NBV agent, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{p}},\mathbf{o}_{t+1})$ is the NBP agent’s local transition.
+The explored transitions are stored in the replay buffer along with demonstration data during training.
 IV-B Viewpoint-centric Voxel Alignment
 […]
-$\mathbf{o}^{\text{v}}=\mathbf{V}\big{(}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}}),\mathbf{e},\mathbf{c}\big{)}$ (3)
+$\mathbf{o}^{\text{i}}=\mathbf{vox}(^{\text{L}}{\text{T}}_{\text{W}}\times\mathbf{o}^{\text{pc}})$ (3)
 […]
-$\mathbf{v}^{*},\mathbf{f}^{*},\mathbf{g}^{*}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}^{*}_{\text{L}},\mathbf{f}^{*}_{\text{L}},\mathbf{g}^{*}_{\text{L}})$ (4)
+[…]
+$\mathbf{v},\mathbf{r},\mathbf{g}=^{\text{W}}{\text{T}}_{\text{L}}\times(\mathbf{v}_{\text{L}}^{*},\mathbf{r}_{\text{L}}^{*},\mathbf{g}_{\text{L}}^{*})$ (4)
 IV-C Viewpoint-aware Demo Augmentation
-To provide early guidance and reduce exploration, we manually alternate camera and gripper movements to generate a small set of demonstration trajectories.
-Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of observation-action pairs collected at a fixed rate.
-To further leverage limited demo data, existing augmentation methods [1, 2, 28] first select key frame indices from the demo trajectory, construct transition samples from the key observation-action pairs, and then resample between key frames to generate additional transitions.
-However, they cannot handle the observation variance caused by viewpoint shifts.
-We define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, identifying key frames for both viewpoint and gripper movements.
-Gripper key frames $k_{j}^{\text{g}}$ are marked when the gripper is stationary or its closure state changes, while key viewpoints are defined as the first stable viewpoint between two gripper key frames, indexed by $k_{j}^{\text{c}}$.
-The key frame indices satisfy $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1.
-As shown in Fig.5, raw transition samples $\text{T}_{j}$ are then constructed as formalized in Eq.(5).
-$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{(\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}),(\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}),r_{j}\}\end{split}$ (5)
-To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$.
-The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
-$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ \end{split}$ (6)
-Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
-Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations(represented by hatched circles) between both viewpoint key frames(solid blue circles) and gripper key frames(solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
+During demonstration, the viewpoint and gripper movements are manually alternated asynchronously, staying stationary while the other is moving.
+These paired stationary viewpoints and gripper positions are called key actions, corresponding to the actions that need to be inferred by the NBV and NBP policies.
+The paired data frame index of these key actions in one demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$, are denoted as $\big{\{}(k^{\text{v}},k^{\text{p}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$ Naively, we can use these key observation-action pairs to construct a prior transition sample set $\{\text{T}\}^{M}$ for active sensor-motor training, where T defined as
+$\text{T}=\{\text{T}^{\text{v}},\text{T}^{\text{p}},r\}=\{(\mathbf{o}_{k^{\text{init}}},\mathbf{a}_{k^{\text{v}}},\mathbf{o}_{k^{\text{v}}}),(\mathbf{o}_{k^{\text{v}}},\mathbf{a}_{k^{\text{p}}},\mathbf{o}_{k^{\text{p}}}),r\}.$ (5)
+However, the number of these raw transition samples is quite limited due to the small demonstration set.
+To fully utilize the demo trajectory data, we uniformly re-sample the initial observations of both $\text{T}^{\text{v}}$ and $\text{T}^{\text{p}}$ to construct additional transition samples for policy training, as shown in Fig.5.
+The resulting additional transitions is formalized as
+$\text{T}_{\text{a}}=\{\text{T}^{\text{v}}_{\text{a}},\text{T}^{\text{p}}_{\text{a}},r\}=\{(\mathbf{o}_{k^{\text{init}}_{\text{a}}},\mathbf{a}_{k^{\text{v}}},\mathbf{o}_{k^{\text{v}}_{\text{a}}}),(\mathbf{o}_{k^{\text{v}}_{\text{a}}},\mathbf{a}_{k^{\text{p}}},\mathbf{o}_{k^{\text{p}}}),r\}.$ (6)
+Both T and $\text{T}_{\text{a}}$ are ultimately stored in the replay buffer as early training experiences for the model.
 IV-D Task-agnostic Auxiliary Reward
-We use sparse reward to avoid designing task-specific rewards.
-However, since sparse reward only provides feedback at the end of an episode, it hinders efficient sensor-motor learning.
-Here we introduce two task-agnostic auxiliary rewards to provide real-time guidance to the training process.
-To encourage interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, $r_{\text{i}}$ provides positive signals when the NBV policy selects a reachable, non-empty ROI and penalizes actions targeting empty or unreachable areas due to execution failures.
-The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
-$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}\end{cases}$ (7)
-To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward($r_{\text{e}}$), inspired by [21].
-The goal is to reduce occupancy entropy within the ROI after viewpoint adjustment, as shown in Fig.6.
-Assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel($e_{i}$) is calculated using the Shannon entropy function(Eq.(8)), where $p_{\text{o}}(i)$ represents its occupancy probability.
-If a voxel is occluded, $p_{\text{o}}$ is set to 0.5.
-$e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
-The total occupancy entropy change of the ROI from time $t$ to $t^{\prime}$ is expressed as:
-$\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
-We define reward $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0.
-The auxiliary reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy.
+Relying solely on a centralized sparse reward to coordinate the collaboration between the NBV and NBP agents may result in only one agent being active while the other remains “lazy”, leading to a non-optimal equilibrium between the two agents [36].
+To mitigate these issues and provide real-time guidance during the training process, we introduce two task-agnostic auxiliary rewards.
+We introduce the Interaction Encouragement Reward $r_{\text{i}}$ to encourage the robot to interact with non-empty ROI areas containing scene entities.
+If the NBP interaction does not fail due to trajectory planning errors or collisions, and the gripper’s finger center point (GCP) is within the ROI after execution, we refer to this ROI as “reachable”.
+The specific value assignment for $r_{\text{i}}$ is detailed as
+$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}.\end{cases}$ (7)
+Figure 5: Demo trajectory augmentation.
+The raw transition T only contains observations and actions at key moments, while the additional augmentative transitions $\text{T}_{\text{a}}$ are constructed from intermediate observations and actions.
+We introduce the Entropy Reduction Reward $r_{\text{e}}$ to encourage NBV agent to select viewpoints that capture more determined visual information of the ROI.
+The reward $r_{\text{e}}$ is defined as $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0 and $\Delta E$ is the total occupancy entropy change of the $N$ voxels inside the ROI after viewpoint adjustment.
+The occupancy entropy of the $i$-th voxel, $e_{i}$, is calculated using the Shannon entropy function [22], shown in Eq.(8), where $p_{\text{o}}(i)$ represents its occupancy probability, and $q_{\text{o}}(i)=1-p_{\text{o}}(i)$.
+$\begin{split}\Delta E&=E_{t}-E_{t^{\prime}}=\frac{1}{N}\big{(}\textstyle\sum_{i=1}^{N}e_{i}^{t}-\textstyle\sum_{i=1}^{N}e_{i}^{t^{\prime}}\big{)}\\ e_{i}&=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-q_{\text{o}}(i)\log_{2}\big{(}q_{\text{o}}(i)\big{)}\end{split}$ (8)
+The reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy.
 Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion.
-These task-agnostic rewards generalize across tasks without the need for manual tuning.
-Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene.
-The model is encouraged to select non-empty ROIs and reduce occluded voxels from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$.
+These rewards are unrelated to specific task objectives, thus generalizing across tasks without the need for manual tuning.
 V Experiment Setup
-Task Setting.
-We trained and evaluated our model on RLBench[9], a widely-used public benchmark for robotic manipulation in reinforcement learning.
-Task selection followed two criteria: 1) each task must include at least one viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across selected tasks.
-As the ROI shifts throughout the manipulation process, and there may be phases without distinct objects(e.g., midair gripper movements), defining occlusion rates based on object segmentation, as done in previous methods[14, 21], is impractical.
-To address this, we define Task Occlusion Rate(TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(10).
-Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation.
-To compare the occlusion levels across selected tasks, we assume that the human-demoed active viewpoint contains the strongest task-specific prior knowledge.
-Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint(oracle).
-$\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (10)
-Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under overhead, front, wrist-mounted cameras as well as the human-demoed active camera.
-Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 100 demonstration trajectories for each task as a benchmark.
+Tasks and Initialization Settings.
+We trained and evaluated our model on RLBench [12], a widely used public benchmark for robotic manipulation.
+The basic scene of RLBench is illustrated in Fig.3, featuring several pre-set cameras, including fixed front and overhead views, as well as a wrist-mounted camera attached to the gripper.
+At the start of each episode, the scene objects are randomly positioned within a legal region of $0.67\text{m}\times 0.91\text{m}$ on the tabletop and randomly rotated around the $z$-axis of the world frame, which is perpendicular to the table.
+This ensures variation in scene initialization, guaranteeing differences in each training and testing episode.
+We selected 8 tasks (Table I) from RLBench following two criteria: 1) each task must include at least one RLBench default viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across the selected tasks.
+To evaluate the occlusion level of tasks and viewpoints, we define the Viewpoint Occlusion Rate (VOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur.
+This is calculated based on observations from that viewpoint across $L=100$ demonstration trajectories, as described in Eq.(9).
+Here, $E^{l,k}$, represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$, after the viewpoint adjustment.
+Furthermore, we normalize the VOR of all viewpoints with reference to the VOR of the human-demonstrated viewpoint in the same task.
+This normalization facilitates intuitive comparisons across tasks and viewpoints.
+The results are shown in Fig.7.
+$\text{VOR}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{k=0}^{K}E^{l,k}$ (9)
+Figure 6: Viewpoint adjustment and ROI selection in a 2D schematic scene.
+We define an ROI as non-empty only if it contains at least one voxel with an occupancy probability of $p_{\text{o}}=1.0$; otherwise, it is considered empty.
 Baselines.
-We compared our model against two baseline categories: 1) visual manipulation model that utilize static third-person and wrist-mounted cameras, and 2) active third-person camera-based models.
-For the first category, we compared our model against C2FARM [2] with static front, overhead and wrist-mounted camera settings to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability.
-For the second category, we compared our model with Behavior Cloning(BC) to test whether it can learn new sensor-motor skills beyond demonstration imitation.
-We also tested a Random viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled(learning rate set to zero).
-This setup allows us to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random choices.
+We compared our model against three baseline categories: 1) visual manipulation models that utilize static third-person and wrist-mounted cameras, 2) active third-person camera-based models, and 3) Oracle baseline, which utilizes fully fused image from all default camera views in RLBench.
+For the first category, we compared our model to C2FARM (C2F) [2] using static front, overhead, wrist and front + wrist (f+w) camera settings to evaluate the effectiveness of our active vision-action model in handling tasks with limited observability.
+For the second category, we evaluated our model against Behavior Cloning (BC) to test its ability to learn new sensor-motor skills beyond demonstration imitation.
+Additionally, we included a Random Viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled (learning rate set to zero).
+This setup assesses the performance gain from active viewpoint adjustment versus random selection.
+For the Oracle baseline, we disabled the NBV viewpoint selection ability since full camera views were always accessible, but retained NBV’s ROI inference capability.
 Training Setting.
-We collected 30 demo trajectories for each task and used them as prior knowledge for training.
-C2FARM was implemented based on its official release and trained with sparse task rewards.
-The Behavior Cloning model was trained with MSE loss, and the Random model used the same reward structure as our model(SectionIV-D).
-All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates, with a batch size of 128.
+We collected 30 demonstration trajectories for each task, for both ours and baseline methods.
+used them as prior knowledge for training.
+The Behavior Cloning model was trained using MSE loss, while the Random model used the same reward as our model but set the NBV learning rate to zero.
+All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates.
 Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
+VI Result
+Main Performance.
+Main Performance.
+We first analyzed the task Success Rate (SR) and Episode Length (EL) across all methods on 8 tasks to evaluate their overall performance.
+Table I shows that our model achieved the highest average success rate, outperforming the second-best single camera setting (C2F+front) by 26%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger.
+Similarly, Table II shows that our model had the shortest average episode length, outperforming the second-best single camera setting (C2F+front) by 28.2% , demonstrating its high proficiency in the tasks.
+Figure 7: Normalized viewpoint occlusion rate of selected tasks. We consider human-demonstrated viewpoints to have the highest task prior knowledge and use them as the baseline for normalization.
+We found that C2F with single camera setting performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, and it occasionally matching or surpassing our method.
+However, their performance dropped sharply in occluded scenarios, leading to a higher performance variance.
+In contrast, our model remained robust in occluded tasks due to its active perception ability.
+Leveraging complementary information from different viewpoints, C2F with a dual-camera setup (front + wrist) outperforms a single camera significantly but remains slightly below our single active camera setup, with both delivering comparable results.
+The Oracle baseline, benefiting from global scene perception, achieved the highest success rate and shortest episode length.
+In addition, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), it struggled with limited training samples, achieving only a 14% success rate.
+Meanwhile the random viewpoint policy performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
 TABLE I: Task success $\uparrow$ . Mean and std of 100 tests are reported (3 seeds).
-Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
-C2FARM(wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
-C2FARM(overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
-C2FARM(front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
+Methods Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
+C2F (wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
+C2F (overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
+C2F (front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
+C2F (f+w) 0.95±0.03 0.32±0.04 0.65±0.04 0.55±0.06 0.84±0.04 0.92±0.02 0.85±0.04 0.73±0.01 0.72±0.23
 Random (active) 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.00±0.00 0.29±0.02 0.00±0.00 0.00±0.00 0.04±0.09
@@
 Ours (active) 0.81±0.02 0.56±0.04 0.73±0.04 0.50±0.03 0.85±0.03 0.96±0.01 0.78±0.02 0.92±0.01 0.76±0.15
+Oracle (full views) 0.95±0.02 0.93±0.03 0.91±0.05 0.82±0.03 0.92±0.01 0.98±0.01 0.97±0.02 0.96±0.04 0.93±0.05
 TABLE II: Task episode $\downarrow$ . The maximum length of each task episode is limited to 10 steps.
-Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
-C2FARM(wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
-C2FARM(overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
-C2FARM(front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
+Methods Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
+C2F (wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
+C2F (overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
+C2F (front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
+C2F (f+w) 3.42±0.21 7.51±0.26 6.14±0.25 6.29±0.35 4.81±0.34 3.13±0.14 5.60±0.17 5.22±0.11 5.27±1.48
 Random (active) 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 10.00±0.00 7.76±0.27 10.00±0.00 10.00±0.00 9.72±0.75
@@
 Ours (active) 5.35±0.08 6.43±0.28 5.15±0.26 6.65±0.19 3.89±0.28 3.33±0.09 4.30±0.17 3.91±0.07 4.88±1.16
-VI Result
-Main Performance.
-We first analyzed the average task success rate(SR) and episode length(EL) across different models to evaluate overall sensor-motor performance.
-Table I shows that our model achieved the highest success rate, outperforming the second-best by 21%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger.
-Similarly, Table II shows that our model had the shortest average manipulation steps, outperforming the second-best by 28.2% and showing the least variance in step counts.
-Fixed and wrist-mounted cameras performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, C2FARM occasionally matching or surpassing our model.
-However, their performance dropped sharply in occluded scenarios, leading to a higher variance.
-In contrast, our model remained robust in occluded tasks.
-Although the BC model demonstrated some ability to adjust its viewpoint(as detailed below), it struggled with limited training samples, achieving only a 14% success rate.
-The random viewpoint model performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
-Figure 8: Partial rollout cases and viewpoint distribution.
-(a) The image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes.
-(b) Red circles indicate manually set viewpoints from 30 demo trajectories, while blue triangles represent the learned viewpoint positions from successful episodes across 100 test runs.
+Oracle (full views) 3.41±0.13 4.56±0.16 4.33±0.38 4.32±0.32 3.49±0.24 3.08±0.01 3.78±0.04 3.12±0.29 3.76±0.58
+Figure 8: Result proportion and partial viewpoint distribution.
+(a) Proportion of success, timeouts, and execution failures across test episodes for different models.
+(b)Red squares indicate manually set viewpoints from 30 demo trajectories, while green circles and blue triangles represent the viewpoints learned by the behavior cloning and our model, respectively, from successful episodes across 100 tests.
 Active Perception Strategy Analysis.
-We further analyzed the perception strategies of all models using two metrics: Average Single Step Information Gain(ASSIG) and Average Initial Information Gain(AIIG).
-ASSIG captures the average reduction in ROI entropy before and after a single viewpoint shift, while AIIG measures the entropy reduction compared to the initial viewpoint.
-$\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (11)
-$\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (12)
-Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the length of episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9).
-The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint.
-As shown in Table III, our model significantly enhances ROI observation by adjusting viewpoints, outperforming the BC model by 2.8% in ASSIG and 1.2% in AIIG.
-In contrast, static front and overhead cameras, which cannot shift viewpoints, fail to achieve information gain.
-The wrist-mounted camera, unable to move independently of the gripper, shows zero ASSIG and is unsuitable for AIIG computation.
-Fig.8(b) compares the distribution of viewpoints learned by our model with the manually set active viewpoints from the demo trajectories used during training.
-Unlike demo patterns, our model’s learned viewpoints are more concentrated in some tasks(Basketball_in_hoop, unplug_charger) and more varied in others(meat_on_grill, open_drawer), indicating it learned to select task-relevant viewpoints rather than simply imitating demo data.
-The camera observations at key moments(Fig.
-8) further showcase our model’s active visual perception capabilities.
-Figure 9: Proportion of success, timeouts, and execution failures across test episodes for different models.
-TABLE III: Comparison of information gain and interaction capacity across different models.
-Models ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
-C2FARM(wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
-C2FARM(overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
-C2FARM(front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
+Active perception involves accurately identifying effective ROI regions in the scene and selecting optimal viewpoints to observe them, thereby supporting task completion.
+We propose Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG) to quantitatively evaluate the model’s ability to enhance ROI observation quality through viewpoint adjustments.
+ASSIG measures single-step information gain, while AIIG quantifies improvement relative to the initial viewpoint.
+They are defined as follows:
+$\text{ASSIG}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{t=0}^{T}\Delta E^{l,t}$ (10)
+$\text{AIIG}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{t=0}^{T}\Delta E^{l,t}_{\text{init}},$ (11)
+Where $L=100$ testing episodes, $T$ is the episode length and $\Delta E$ are defined in Eq.(8).
+Fig.8(b) shows that both our model and the BC model demonstrate some learned viewpoint adjustment ability.
+Table.III further confirms this, with our model achieving up to 4.6% and 7.0% higher ASSIG and AIIG scores than the BC model.
+However, as Fig.8(a) shows, the BC model has significantly more failures due to timeouts, indicating difficulty in accurately localizing ROI regions from limited demos.
+Additionally, our model learns more task-relevant viewpoints: some tasks (e.g., basketball_in_hoop, put_rubbish_in_bin) show more focused distributions, while others (e.g., meat_on_grill, open_drawer) exhibit greater generalization, reflecting further skill acquisition through interaction rather than mere imitation as seen with BC.
+Environmental Modeling and Interactive Capabilities.
+We evaluated the models’ environment modeling and interaction capabilities by recording Success, Timeout (TO), and Execution Failure (EF) rates over 100 test episodes (Fig.8).
+The C2F model exhibited a high failure rate (up to 42%) due to trajectory execution failures.
+Similarly, the Random model saw the majority of its test episodes end in failure because of execution errors.
+Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts.
+In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
+As shown in Table III, we also analyzed the average steps to First Interaction (FI) and Non-interactive steps (NI) per model.
+The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise ROI location and action skills from limited demonstrations.
+In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
+TABLE III: Comparison of information gain and interaction capacity across different methods.
+Methods ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
+C2F (wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
+C2F (overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
+C2F (front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
+C2F (f+w) 0.0/0.0 N/A 1.91±0.55 2.63±1.32
 Random (active) 0.0/5.1% 0.0/5.2% 9.07±1.47 9.49±1.07
@@
 Ours (active) 3.5%/6.9% 6.8%/25.1% 1.64±0.61 2.15±1.03
-Environmental Modeling and Interactive Capabilities.
-We evaluated the models’ environment modeling and interaction capabilities by recording success, timeout(TO), and execution failure(EF) rates over 100 test episodes(Fig.9).
-The C2FARM model exhibited a high failure rate(up to 42%) due to trajectory execution failures.
-Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors.
-Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts.
-In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
-As shown in Table III, we also analyzed the average steps to first interaction(FI) and non-interactive steps(NI) per model.
-The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise manipulations from limited demonstrations.
-In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
+Oracle (full views) 0.0%/0.0% N/A 1.58±1.36 1.86±1.56
 Ablation Study.
-We evaluated the impact of three strategies on manipulation performance:(1) viewpoint-centric voxel alignment(Section IV-B), (2) viewpoint-aware demo augmentation(Section IV-C), and (3) task-agnostic auxiliary rewards(Section IV-D).
-Separate ablation tests were conducted for each, as well as a combined test removing all strategies.
+We first evaluated the impact of three strategies on manipulation performance: viewpoint-centric voxel alignment (Sec IV.B), viewpoint-aware demo augmentation (Sec IV.C), and task-agnostic auxiliary rewards (Sec IV.D).
 Table IV shows that each strategy effectively improved success rates (SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
@@
 Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.
+We also evaluated our model with reduced demostrations.
+As shown in Table., reducing the number of demos from 30 to 15 resulted in a 21% drop in success rate.
+However, even with only 15 demos, the model slightly outperformed C2F with a single-camera setup trained on 30 demos, demonstrating its efficiency in leveraging limited demonstrations.
+Without any demos, the success rate dropped to 8%, highlighting the significant challenges of learning active vision-action skills in a large observation-action space.
+Reducing the number of demos also increased episode lengths and execution failures, emphasizing the importance of providing sufficient, high-quality samples early in training.
+Such samples can help the model quickly learn environmental geometry and facilitate more efficient action exploration in subsequent stages.
+Interestingly, we observed that the model achieved meaningful ASSIG scores even without any demo data, slightly exceeding those obtained with 15 demos.
+We speculate that this is because when the global agent struggled to discover sparse task rewards, the local NBV agent may focused more on short-term entropy rewards $r_{\text{e}}$, leading to higher ASSIG scores.
+However, the AIIG scores, which require long-term planning, remained lower than those achieved with 15 demos.
 TABLE IV: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards.
 Ablation Setting Result Proportion Interactivity Sensing
-align aug aux SR EF TO EL FI NI AIIG ASSIG
+[…]
 0.34 0.53 0.14 7.61 4.33 5.81 0.046 0.082
@@
 0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035
+TABLE V: Ablation over different demo quantities.
+Demo Result Proportion Interactivity Sensing
+[…]
+0 0.08 0.70 0.22 9.48 7.99 8.76 0.019 0.026
+15 0.51 0.36 0.13 6.53 3.23 4.38 0.046 0.020
+30 0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035
 VII Conclusion and Future Work
 In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model.
-Our model first infers the optimal viewpoint based on the task goal, then selects the manipulation action from the updated observation.
-To enhance performance and training stability, we integrate voxel alignment for spatial consistency, demonstration augmentation for enhanced learning, and auxiliary rewards for real-time training guidance.
-These strategies enable the model to handle occlusions and improve sensor-motor coordination.
-Experimental results show that our model significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment.
-However, complex occlusions and dynamic environments remain challenging.
-Future work could explore real-time visual feedback to better adapt to environmental changes.
+Our model first infers the optimal viewpoint to actively observe the scene, then selects the manipulation action from the updated observation.
+Our model was trained and evaluated on eight RLBench tasks with varying levels of occlusion, demonstrating that it significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment.
+However, one limitation is its reliance on a small number of demonstration samples to provide early training signals.
+Future work could explore strategies for zero-shot training and extend the approach to longer-horizon tasks.